Using as

The GNU Assembler

March 1993

The Free Software Foundation Inc. thanks The Nice Computer Company of Australia for loaning Dean Elsner to write the first (Vax) version of as for Project GNU. The proprietors, management and staff of TNCCA thank FSF for distracting the boss while they got some work done.

Dean Elsner, Jay Fenlason & friends

Using as

Edited by Roland Pesch for Cygnus Support

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991, 1992, 1993 Free Software Foundation, Inc.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided also that the section entitled “GNU General Public License” is included exactly as in the original, and provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that the section entitled “GNU General Public License” may be included in a translation approved by the Free Software Foundation instead of in the original English.

1 Overview

This manual is a user guide to the GNU assembler as.

Here is a brief summary of how to invoke as. For details, see Chapter 2 [Comand-Line Options], page 7.

       as [ -a[dhlns] ] [ -D ] [ -f ]
        [ -I path ] [ -K ] [ -L ]
        [ -o objfile ] [ -R ] [ -v ] [ -w ]
        [ -Av6 | -Av7 | -Av8 | -Asparclite | -bump ]
        [ -ACA | -ACA_A | -ACB | -ACC | -AKA | -AKB | -AKC | -AMC ]
        [ -b ] [ -norelax ]
        [ -l ] [ -m68000 | -m68010 | -m68020 | ... ]

[ -- | files …]

-a[dhlns]

Turn on listings; ‘-ad’, omit debugging pseudo-ops from listing, ‘-ah’, include highlevel source, ‘-al’, assembly listing, ‘-an’, no forms processing, ‘-as’, symbols. These options may be combined; e.g., ‘-aln’ for assembly listing without forms processing. By itself, ‘-a’ defaults to ‘-ahls’ — that is, all listings turned on.

-D This option is accepted only for script compatibility with calls to other assemblers; it

has no effect on as.

-f “fast”—skip preprocessing (assume source is compiler output) -I path Add path to the search list for .include directives -K Issue warnings when difference tables altered for long displacements. -L Keep (in symbol table) local symbols, starting with ‘L’ -o objfile Name the object-file output from as -R Fold data section into text section -v Announce as version -W Suppress warning messages -- | files …

Standard input, or source files to assemble.

The following options are available when as is configured for the Intel 80960 processor.

-ACA | -ACA_A | -ACB | -ACC | -AKA | -AKB | -AKC | -AMC

Specify which variant of the 960 architecture is the target.

-b Add code to collect statistics about branches taken. -norelax Do not alter compare-and-branch instructions for long displacements; error if necessary.

The following options are available when as is configured for the Motorola 68000 series.

-l Shorten references to undefined symbols, to one word instead of two.

-m68000 | -m68008 | -m68010 | -m68020 | -m68030 | -m68040
| -m68302 | -m68331 | -m68332 | -m68333 | -m68340 | -mcpu32

Specify what processor in the 68000 family is the target. The default is normally the 68020, but this can be changed at configuration time.

-m68881 | -m68882 | -mno-68881 | -mno-68882

The target machine does (or does not) have a floating-point coprocessor. The default is to assume a coprocessor for 68020, 68030, and cpu32. Although the basic 68000 is not compatible with the 68881, a combination of the two can be specified, since it’s possible to do emulation of the coprocessor instructions with the main processor.

-m68851 | -mno-68851

The target machine does (or does not) have a memory-management unit coprocessor. The default is to assume an MMU for 68020 and up.

The following options are available when as is configured for the SPARC architecture:

-Av6 | -Av7 | -Av8 | -Asparclite

Explicitly select a variant of the SPARC architecture.

-bump Warn when the assembler switches to another architecture.

1.1 Structure of this Manual

This manual is intended to describe what you need to know to use gnu as. We cover the syntax expected in source files, including notation for symbols, constants, and expressions; the directives that as understands; and of course how to invoke as.

This manual also describes some of the machine-dependent features of various flavors of the assembler.

On the other hand, this manual is not intended as an introduction to programming in assembly language—let alone programming in general! In a similar vein, we make no attempt to introduce the machine architecture; we do not describe the instruction set, standard mnemonics, registers or addressing modes that are standard to a particular architecture. You may want to consult the manufacturer’s machine architecture manual for this information.

1.2 as, the GNU Assembler

GNU as is really a family of assemblers. If you use (or have used) the GNU assembler on one architecture, you should find a fairly similar environment when you use it on another architecture. Each version has much in common with the others, including object file formats, most assembler directives (often called pseudo-ops) and assembler syntax.

as is primarily intended to assemble the output of the GNU C compiler gcc for use by the linker ld. Nevertheless, we’ve tried to make as assemble correctly everything that other assemblers for the same machine would assemble. Any exceptions are documented explicitly (see Chapter 8 [Machine Dependencies], page 49). This doesn’t mean as always uses the same syntax as another assembler for the same architecture; for example, we know of several incompatible versions of 680x0 assembly language syntax.

Unlike older assemblers, as is designed to assemble a source program in one pass of the source file. This has a subtle impact on the .org directive (see Section 7.37 [.org], page 41).

1.3 Object File Formats

The GNU assembler can be configured to produce several alternative object file formats. For the most part, this does not affect how you write assembly language programs; but directives for debugging symbols are typically different in different file formats. See Section 5.5 [Symbol Attributes], page 27.

1.4 Command Line

After the program name as, the command line may contain options and file names. Options may appear in any order, and may be before, after, or between file names. The order of file names is significant.

‘--’ (two hyphens) by itself names the standard input file explicitly, as one of the files for as to assemble.

Except for ‘--’ any command line argument that begins with a hyphen (‘-’) is an option. Each option changes the behavior of as. No option changes the way another option works. An option is a ‘-’ followed by one or more letters; the case of the letter is important. All options are optional.

Some options expect exactly one file name to follow them. The file name may either immediately follow the option’s letter (compatible with older assemblers) or it may be the next command argument (GNU standard). These two command lines are equivalent:

     as -o my-object-file.o mumble.s
     as -omy-object-file.o mumble.s

1.5 Input Files

We use the phrase source program, abbreviated source, to describe the program input to one run of as. The program may be in one or more files; how the source is partitioned into files doesn’t change the meaning of the source.

The source program is a concatenation of the text in all the files, in the order specified.

Each time you run as it assembles exactly one source program. The source program is made up of one or more files. (The standard input is also a file.)

You give as a command line that has zero or more input file names. The input files are read (from left file name to right). A command line argument (in any position) that has no special meaning is taken to be an input file name.

If you give as no file names it attempts to read one input file from the as standard input, which is normally your terminal. You may have to type CTL-D to tell as there is no more program to assemble.

Use ‘--’ if you need to explicitly name the standard input file in your command line.

If the source is empty, as will produce a small, empty object file.

Filenames and Line-numbers

There are two ways of locating a line in the input file (or files) and either may be used in reporting error messages. One way refers to a line number in a physical file; the other refers to a line number in a “logical” file. See Section 1.7 [Error and Warning Messages], page 6.

Physical files are those files named in the command line given to as.

Logical files are simply names declared explicitly by assembler directives; they bear no relation to physical files. Logical file names help error messages reflect the original source file, when as source is itself synthesized from other files. See Section 7.4 [.app-file], page 34.

1.6 Output (Object) File

Every time you run as it produces an output file, which is your assembly language program translated into numbers. This file is the object file, named b.out, if as is configured for the Intel 80960, or unless you tell as to give it another name by using the -o option. Conventionally, object file names end with ‘.o’. The default name of ‘a.out’ is used for historical reasons: older assemblers were capable of assembling self-contained programs directly into a runnable program. (For some formats, this isn’t currently possible, but it can be done for a.out format.)

The object file is meant for input to the linker ld. It contains assembled program code, information to help ld integrate the assembled program into a runnable file, and (optionally) symbolic information for the debugger.

1.7 Error and Warning Messages

as may write warnings and error messages to the standard error file (usually your terminal). This should not happen when a compiler runs as automatically. Warnings report an assumption made so that as could keep assembling a flawed program; errors report a grave problem that stops the assembly.

Warning messages have the format

     file_name:NNN:Warning Message Text

(where NNN is a line number). If a logical file name has been given (see Section 7.4 [.app-file], page 34) it is used for the filename, otherwise the name of the current input file is used. If a logical line number was given (see Section 7.31 [.line], page 40) then it is used to calculate the number printed, otherwise the actual line in the current source file is printed. The message text is intended to be self explanatory (in the grand Unix tradition).

Error messages have the format

     file_name:NNN:FATAL:Error Message Text

The file name and line number are derived as for warning messages. The actual message text may be rather less explanatory because many of them aren’t supposed to happen.

2 Command-Line Options

This chapter describes command-line options available in all versions of the GNU assembler; see Chapter 8 [Machine Dependencies], page 49, for options specific to particular machine architectures.

If you are invoking as via the GNU C compiler (version 2), you can use the ‘-Wa’ option to pass arguments through to the assembler. The assembler arguments must be separated from each other (and the ‘-Wa’) by commas. For example:

     gcc -c -g -O -Wa,-alh,-L file.c

will cause a listing to be emitted to standard output with high-level and assembly source.

Many compiler command-line options, such as ‘-R’ and many machine-specific options, will be automatically be passed to the assembler by the compiler, so usually you do not need to use this ‘-Wa’ mechanism.

2.1 Enable Listings: -a[dhlns]

These options enable listing output from the assembler. By itself, ‘-a’ requests high-level, assembly, and symbols listing. Other letters may be used to select specific options for the list: ‘-ah’ requests a high-level language listing, ‘-al’ requests an output-program assembly listing, and ‘-as’ requests a symbol table listing. High-level listings require that a compiler debugging option like ‘-g’ be used, and that assembly listings (‘-al’) be requested also.

The ‘-ad’ option may be used to omit debugging pseudo-ops from the listing.

Once you have specified one of these options, you can further control listing output and its appearance using the directives .list, .nolist, .psize, .eject, .title, and .sbttl. The ‘-an’ option turns off all forms processing. If you do not request listing output with one of the ‘-a’ options, the listing-control directives have no effect.

The letters after ‘-a’ may be combined into one option, e.g., ‘-aln’.

2.2 -D

This option has no effect whatsoever, but it is accepted to make it more likely that scripts written for other assemblers will also work with as.

2.3 Work Faster: -f

‘-f’ should only be used when assembling programs written by a (trusted) compiler. ‘-f’ stops the assembler from pre-processing the input file(s) before assembling them. See Section 3.1 [Preprocessing], page 11.

Warning: if the files actually need to be pre-processed (if they contain comments, for example), as will not work correctly if ‘-f’ is used.

2.4 .include search path: -I path

Use this option to add a path to the list of directories as will search for files specified in .include directives (see Section 7.27 [.include], page 39). You may use -I as many times as necessary to include a variety of paths. The current working directory is always searched first; after that, as searches any ‘-I’ directories in the same order as they were specified (left to right) on the command line.

2.5 Difference Tables: -K

as sometimes alters the code emitted for directives of the form ‘.word sym1-sym2’; see Section 7.54 [.word], page 46. You can use the ‘-K’ option if you want a warning issued when this is done.

2.6 Include Local Labels: -L

Labels beginning with ‘L’ (upper case only) are called local labels. See Section 5.3 [Symbol Names], page 25. Normally you don’t see such labels when debugging, because they are intended for the use of programs (like compilers) that compose assembler programs, not for your notice. Normally both as and ld discard such labels, so you don’t normally debug with them.

This option tells as to retain those ‘L…’ symbols in the object file. Usually if you do this you also tell the linker ld to preserve symbols whose names begin with ‘L’.

2.7 Name the Object File: -o

There is always one object file output when you run as. By default it has the name ‘a.out’ (or ‘b.out’, for Intel 960 targets only). You use this option (which takes exactly one filename) to give the object file a different name.

Whatever the object file is called, as will overwrite any existing file of the same name.

2.8 Join Data and Text Sections: -R

-R tells as to write the object file as if all data-section data lives in the text section. This is only done at the very last moment: your binary data are the same, but data section parts are relocated differently. The data section part of your object file is zero bytes long because all its bytes are appended to the text section. (See Chapter 4 [Sections and Relocation], page 19.)

When you specify -R it would be possible to generate shorter address displacements (because we don’t have to cross between text and data section). We refrain from doing this simply for compatibility with older versions of as. In future, -R may work this way.

When as is configured for COFF output, this option is only useful if you use sections named

‘.text’ and ‘.data’.

2.9 Announce Version: -v

You can find out what version of as is running by including the option ‘-v’ (which you can also spell as ‘-version’) on the command line.

2.10 Suppress Warnings: -W

as should never give a warning or error message when assembling compiler output. But programs written by people often cause as to give a warning that a particular assumption was made. All such warnings are directed to the standard error file. If you use this option, no warnings are issued. This option only affects the warning messages: it does not change any particular of how as assembles your file. Errors, which stop the assembly, are still reported.

3 Syntax

This chapter describes the machine-independent syntax allowed in a source file. as syntax is similar to what many other assemblers use; it is inspired by the BSD 4.2 assembler, except that as does not assemble Vax bit-fields.

3.1 Pre-Processing

The pre-processor:

• adjusts and removes extra whitespace. It leaves one space or tab before the keywords on a line, and turns any other whitespace on the line into a single space.

• removes all comments, replacing them with a single space, or an appropriate number of newlines.

• converts character constants into the appropriate numeric values.

Excess whitespace, comments, and character constants cannot be used in the portions of the input text that are not pre-processed.

If the first line of an input file is #NO_APP or the ‘-f’ option is given, the input file will not be pre-processed. Within such an input file, parts of the file can be pre-processed by putting a line that says #APP before the text that should be pre-processed, and putting a line that says #NO_APP after them. This feature is mainly intend to support asm statements in compilers whose output normally does not need to be pre-processed.

3.2 Whitespace

Whitespace is one or more blanks or tabs, in any order. Whitespace is used to separate symbols, and to make programs neater for people to read. Unless within character constants (see Section 3.6.1 [Character Constants], page 14), any whitespace means the same as exactly one space.

3.3 Comments

There are two ways of rendering comments to as. In both cases the comment is equivalent to one space.

Anything from ‘/*’ through the next ‘*/’ is a comment. This means you may not nest these comments.

     /*
       The only way to include a newline ('\n') in a comment
       is to use this sort of comment.
     */
     /* This sort of comment does not nest. */

Anything from the line comment character to the next newline is considered a comment and is ignored. The line comment character is ‘#’ on the Vax; ‘#’ on the i960; ‘!’ on the SPARC; ‘|’ on the 680x0; ‘;’ for the AMD 29K family; ‘;’ for the H8/300 family; ‘!’ for the H8/500 family; ‘!’ for the Hitachi SH; ‘!’ for the Z8000; see Chapter 8 [Machine Dependencies], page 49.

On some machines there are two different line comment characters. One will only begin a comment if it is the first non-whitespace character on a line, while the other will always begin a comment.

To be compatible with past assemblers, a special interpretation is given to lines that begin with ‘#’. Following the ‘#’ an absolute expression (see Chapter 6 [Expressions], page 29) is expected: this will be the logical line number of the next line. Then a string (See Section 3.6.1.1 [Strings], page 14.) is allowed: if present it is a new logical file name. The rest of the line, if any, should be whitespace.

If the first non-whitespace characters on the line are not numeric, the line is ignored. (Just like a comment.)

                               # This is an ordinary comment.
     # 42-6 "new_file_name"    # New logical file name
                               # This is logical line # 36.

This feature is deprecated, and may disappear from future versions of as.

3.4 Symbols

A symbol is one or more characters chosen from the set of all letters (both upper and lower case), digits and the three characters ‘_.$’. On most machines, you can also use $ in symbol names; exceptions are noted in Chapter 8 [Machine Dependencies], page 49. No symbol may begin with a digit. Case is significant. There is no length limit: all characters are significant. Symbols are delimited by characters not in that set, or by the beginning of a file (since the source program must end with a newline, the end of a file is not a possible symbol delimiter). See Chapter 5 [Symbols], page 25.

3.5 Statements

A statement ends at a newline character (‘\n’) or line separator character. (The line separator is usually ‘;’, unless this conflicts with the comment character; see Chapter 8 [Machine Dependencies], page 49.) The newline or separator character is considered part of the preceding statement. Newlines and separators within character constants are an exception: they don’t end statements.

It is an error to end any statement with end-of-file: the last character of any input file should be a newline.

You may write a statement on more than one line if you put a backslash (\) immediately in front of any newlines within the statement. When as reads a backslashed newline both characters are ignored. You can even put backslashed newlines in the middle of symbol names without changing the meaning of your source program.

An empty statement is allowed, and may include whitespace. It is ignored.

A statement begins with zero or more labels, optionally followed by a key symbol which determines what kind of statement it is. The key symbol determines the syntax of the rest of the statement. If the symbol begins with a dot ‘.’ then the statement is an assembler directive: typically valid for any computer. If the symbol begins with a letter the statement is an assembly language instruction: it will assemble into a machine language instruction. Different versions of as for different computers will recognize different instructions. In fact, the same symbol may represent a different instruction in a different computer’s assembly language.

A label is a symbol immediately followed by a colon (:). Whitespace before a label or after a colon is permitted, but you may not have whitespace between a label’s symbol and its colon. See Section 5.1 [Labels], page 25.

     label:     .directive    followed by something
     another_label:           # This is an empty statement.
                instruction   operand_1, operand_2, . . .

3.6 Constants

A constant is a number, written so that its value is known by inspection, without knowing any context. Like this:

     .byte  74, 0112, 092, 0x4A, 0X4a, 'J, '\J # All the same value.
     .ascii "Ring the bell\7"                  # A string constant.
     .octa  0x123456789abcdef0123456789ABCDEF0 # A bignum.
     .float 0f-314159265358979323846264338327\
     95028841971.693993751E-40                 # - pi, a flonum.

3.6.1 Character Constants

There are two kinds of character constants. A character stands for one character in one byte and its value may be used in numeric expressions. String constants (properly called string literals) are potentially many bytes and their values may not be used in arithmetic expressions.

3.6.1.1 Strings

A string is written between double-quotes. It may contain double-quotes or null characters. The way to get special characters into a string is to escape these characters: precede them with a backslash ‘\’ character. For example ‘\\’ represents one backslash: the first \ is an escape which tells as to interpret the second character literally as a backslash (which prevents as from recognizing the second \ as an escape character). The complete list of escapes follows.

\b Mnemonic for backspace; for ASCII this is octal code 010. \f Mnemonic for FormFeed; for ASCII this is octal code 014. \n Mnemonic for newline; for ASCII this is octal code 012. \r Mnemonic for carriage-Return; for ASCII this is octal code 015. \t Mnemonic for horizontal Tab; for ASCII this is octal code 011.

\ digit digit digit

An octal character code. The numeric code is 3 octal digits. For compatibility with other Unix systems, 8 and 9 are accepted as digits: for example, \008 has the value 010, and \009 the value 011.

\\ Represents one ‘\’ character.

\" Represents one ‘"’ character. Needed in strings to represent this character, because an unescaped ‘"’ would end the string.

\ anything-else

Any other character when escaped by \ will give a warning, but assemble as if the ‘\’ was not present. The idea is that if you used an escape sequence you clearly didn’t want the literal interpretation of the following character. However as has no other interpretation, so as knows it is giving you the wrong code and warns you of the fact.

Which characters are escapable, and what those escapes represent, varies widely among assemblers. The current set is what we think the BSD 4.2 assembler recognizes, and is a subset of what most C compilers recognize. If you are in doubt, don’t use an escape sequence.

3.6.1.2 Characters

A single character may be written as a single quote immediately followed by that character. The same escapes apply to characters as to strings. So if you want to write the character backslash, you must write '\\ where the first \ escapes the second \. As you can see, the quote is an acute accent, not a grave accent. A newline immediately following an acute accent is taken as a literal character and does not count as the end of a statement. The value of a character constant in a numeric expression is the machine’s byte-wide code for that character. as assumes your character code is ASCII: 'A means 65, 'B means 66, and so on.

3.6.2 Number Constants

as distinguishes three kinds of numbers according to how they are stored in the target machine. Integers are numbers that would fit into an int in the C language. Bignums are integers, but they are stored in more than 32 bits. Flonums are floating point numbers, described below.

3.6.2.1 Integers

A binary integer is ‘0b’ or ‘0B’ followed by zero or more of the binary digits ‘01’.

An octal integer is ‘0’ followed by zero or more of the octal digits (‘01234567’).

A decimal integer starts with a non-zero digit followed by zero or more digits (‘0123456789’). A hexadecimal integer is ‘0x’ or ‘0X’ followed by one or more hexadecimal digits chosen from

‘0123456789abcdefABCDEF’.

Integers have the usual values. To denote a negative integer, use the prefix operator ‘-’ discussed under expressions (see Section 6.2.3 [Prefix Operators], page 30).

3.6.2.2 Bignums

A bignum has the same syntax and semantics as an integer except that the number (or its negative) takes more than 32 bits to represent in binary. The distinction is made because in some places integers are permitted while bignums are not.

3.6.2.3 Flonums

A flonum represents a floating point number. The translation is indirect: a decimal floating point number from the text is converted by as to a generic binary floating point number of more than sufficient precision. This generic floating point number is converted to a particular computer’s floating point format (or formats) by a portion of as specialized to that computer.

A flonum is written by writing (in order)

• The digit ‘0’.

• A letter, to tell as the rest of the number is a flonum. e is recommended. Case is not important. On the H8/300, H8/500, Hitachi SH, and AMD 29K architectures, the letter must be one of the letters ‘DFPRSX’ (in upper or lower case). On the Intel 960 architecture, the letter must be one of the letters ‘DFT’ (in upper or lower case).

• An optional sign: either ‘+’ or ‘-’.

• An optional integer part: zero or more decimal digits.

• An optional fractional part: ‘.’ followed by zero or more decimal digits.

• An optional exponent, consisting of:

• An ‘E’ or ‘e’.

• Optional sign: either ‘+’ or ‘-’.

• One or more decimal digits.

At least one of the integer part or the fractional part must be present. The floating point number has the usual base-10 value.

as does all processing using integers. Flonums are computed independently of any floating point hardware in the computer running as.

18 Using as

4 Sections and Relocation

4.1 Background

Roughly, a section is a range of addresses, with no gaps; all data “in” those addresses is treated the same for some particular purpose. For example there may be a “read only” section.

The linker ld reads many object files (partial programs) and combines their contents to form a runnable program. When as emits an object file, the partial program is assumed to start at address 0. ld will assign the final addresses the partial program occupies, so that different partial programs don’t overlap. This is actually an over-simplification, but it will suffice to explain how as uses sections.

ld moves blocks of bytes of your program to their run-time addresses. These blocks slide to their run-time addresses as rigid units; their length does not change and neither does the order of bytes within them. Such a rigid unit is called a section. Assigning run-time addresses to sections is called relocation. It includes the task of adjusting mentions of object-file addresses so they refer to the proper run-time addresses. For the H8/300 and H8/500, and for the Hitachi SH, as pads sections if needed to ensure they end on a word (sixteen bit) boundary.

An object file written by as has at least three sections, any of which may be empty. These are named text, data and bss sections.

When it generates COFF output, as can also generate whatever other named sections you specify using the ‘.section’ directive (see Section 7.42 [.section], page 43). If you don’t use any directives that place output in the ‘.text’ or ‘.data’ sections, these sections will still exist, but will be empty.

Within the object file, the text section starts at address 0, the data section follows, and the bss section follows the data section.

To let ld know which data will change when the sections are relocated, and how to change that data, as also writes to the object file details of the relocation needed. To perform relocation ld must know, each time an address in the object file is mentioned:

• Where in the object file is the beginning of this reference to an address?

• How long (in bytes) is this reference?

• Which section does the address refer to? What is the numeric value of

(address) − (start-address of section)?

• Is the reference to an address “Program-Counter relative”?

In fact, every address as ever uses is expressed as

(section) + (offset into section)

Further, every expression as computes is of this section-relative nature. Absolute expression means an expression with section “absolute” (see Section 4.2 [Ld Sections], page 21). A pass1 expression means an expression with section “pass1” (see Section 4.3 [as Internal Sections], page 22). In this manual we use the notation {secname N} to mean “offset N into section secname”.

Apart from text, data and bss sections you need to know about the absolute section. When ld mixes partial programs, addresses in the absolute section remain unchanged. For example, address {absolute 0} is “relocated” to run-time address 0 by ld. Although two partial programs’ data sections will not overlap addresses after linking, by definition their absolute sections will overlap. Address {absolute 239} in one partial program will always be the same address when the program is running as address {absolute 239} in any other partial program.

The idea of sections is extended to the undefined section. Any address whose section is unknown at assembly time is by definition rendered {undefined U}—where U will be filled in later. Since numbers are always defined, the only way to generate an undefined address is to mention an undefined symbol. A reference to a named common block would be such a symbol: its value is unknown at assembly time so it has section undefined.

By analogy the word section is used to describe groups of sections in the linked program. ld puts all partial programs’ text sections in contiguous addresses in the linked program. It is customary to refer to the text section of a program, meaning all the addresses of all partial program’s text sections. Likewise for data and bss sections.

Some sections are manipulated by ld; others are invented for use of as and have no meaning except during assembly.

4.2 ld Sections

ld deals with just four kinds of sections, summarized below.

named sections text section data section

These sections hold your program. as and ld treat them as separate but equal sections. Anything you can say of one section is true another. When the program is running, however, it is customary for the text section to be unalterable. The text section is often shared among processes: it will contain instructions, constants and the like. The data section of a running program is usually alterable: for example, C variables would be stored in the data section.

bss section This section contains zeroed bytes when your program begins running. It is used to hold unitialized variables or common storage. The length of each partial program’s bss section is important, but because it starts out containing zeroed bytes there is no need to store explicit zero bytes in the object file. The bss section was invented to eliminate those explicit zeros from object files.

absolute section Address 0 of this section is always “relocated” to runtime address 0. This is useful if you want to refer to an address that ld must not change when relocating. In this sense we speak of absolute addresses being “unrelocatable”: they don’t change during relocation.

undefined section This “section” is a catch-all for address references to objects not in the preceding sections.

An idealized example of three relocatable sections follows. The example uses the traditional section names ‘.text’ and ‘.data’. Memory addresses are on the horizontal axis. Partial program #1:

text          data       bss
ttttt        dddd      00

Partial program #2:

text  data     bss
TTT    DDDD    000

linked program:

text                     data                bss
TTT     ttttt             dddd      DDDD     00000  . . .

addresses: 0…

4.3 as Internal Sections

These sections are meant only for the internal use of as. They have no meaning at run-time. You don’t really need to know about these sections for most purposes; but they can be mentioned in as warning messages, so it might be helpful to have an idea of their meanings to as. These sections are used to permit the value of every expression in your assembly language program to be a section-relative address.

absent An expression was expected and none was found. ASSEMBLER-INTERNAL-LOGIC-ERROR!

An internal assembler logic error has been found. This means there is a bug in the assembler.

bignum/flonum If a number can’t be written as a C int constant (a bignum or a flonum, but not an integer), it is recorded as belonging to this “section”. as has to remember that a flonum or a bignum does not fit into 32 bits, and cannot be an argument (see Section 6.2.1 [Arguments], page 29) in an expression: this is done by making a flonum or bignum be in a separate internal section. This is purely for internal as convenience; bignum/flonum section behaves similarly to absolute section.

pass1 section The expression was impossible to evaluate in the first pass. The assembler will attempt a second pass (second reading of the source) to evaluate the expression. Your expression mentioned an undefined symbol in a way that defies the one-pass (section + offset in section) assembly process. No compiler need emit such an expression.

Warning: the second pass is currently not implemented. as will abort with an error message if one is required.

difference section

As an assist to the C compiler, expressions of the forms (undefined symbol) − (expression) something − (undefined symbol) (undefined symbol) − (undefined symbol) are permitted, and belong to the difference section. as re-evaluates such expressions after the source file has been read and the symbol table built. If by that time there are no undefined symbols in the expression then the expression assumes a new section. The intention is to permit statements like ‘.word label - base_of_table’ to be assembled in one pass where both label and base_of_table are undefined. This is useful for compiling C and Algol switch statements, Pascal case statements, FORTRAN computed goto statements and the like.

4.4 Sub-Sections

Assembled bytes conventionally fall into two sections: text and data. You may have separate groups of data in named sections that you want to end up near to each other in the object file, even though they are not contiguous in the assembler source. as allows you to use subsections for this purpose. Within each section, there can be numbered subsections with values from 0 to 8192. Objects assembled into the same subsection will be grouped with other objects in the same subsection when they are all put into the object file. For example, a compiler might want to store constants in the text section, but might not want to have them interspersed with the program being assembled. In this case, the compiler could issue a ‘.text 0’ before each section of code being output, and a ‘.text 1’ before each group of constants being output.

Subsections are optional. If you don’t use subsections, everything will be stored in subsection number zero.

Each subsection is zero-padded up to a multiple of four bytes. (Subsections may be padded a different amount on different flavors of as.)

Subsections appear in your object file in numeric order, lowest numbered to highest. (All this to be compatible with other people’s assemblers.) The object file contains no representation of subsections; ld and other programs that manipulate object files will see no trace of them. They just see all your text subsections as a text section, and all your data subsections as a data section.

To specify which subsection you want subsequent statements assembled into, use a numeric argument to specify it, in a ‘.text expression’ or a ‘.data expression’ statement. When generating COFF output, you can also use an extra subsection argument with arbitrary named sections: ‘.section name, expression’. Expression should be an absolute expression. (See Chapter 6 [Expressions], page 29.) If you just say ‘.text’ then ‘.text 0’ is assumed. Likewise ‘.data’ means ‘.data 0’. Assembly begins in text 0. For instance:

     .text 0     # The default subsection is text 0 anyway.
     .ascii "This lives in the first text subsection. *"
     .text 1
     .ascii "But this lives in the second text subsection."
     .data 0
     .ascii "This lives in the data section,"
     .ascii "in the first data subsection."
     .text 0
     .ascii "This lives in the first text section,"
     .ascii "immediately following the asterisk (*)."

Each section has a location counter incremented by one for every byte assembled into that section. Because subsections are merely a convenience restricted to as there is no concept of a subsection location counter. There is no way to directly manipulate a location counter—but the .align directive will change it, and any label definition will capture its current value. The location counter of the section that statements are being assembled into is said to be the active location counter.

4.5 bss Section

The bss section is used for local common variable storage. You may allocate address space in the bss section, but you may not dictate data to load into it before your program executes. When your program starts running, all the contents of the bss section are zeroed bytes.

Addresses in the bss section are allocated with special directives; you may not assemble anything directly into the bss section. Hence there are no bss subsections. See Section 7.8 [.comm], page 34, see Section 7.29 [.lcomm], page 39.

5 Symbols

Symbols are a central concept: the programmer uses symbols to name things, the linker uses symbols to link, and the debugger uses symbols to debug.

Warning: as does not place symbols in the object file in the same order they were declared. This may break some debuggers.

5.1 Labels

A label is written as a symbol immediately followed by a colon ‘:’. The symbol then represents the current value of the active location counter, and is, for example, a suitable instruction operand. You are warned if you use the same symbol to represent two different locations: the first definition overrides any other definitions.

5.2 Giving Symbols Other Values

A symbol can be given an arbitrary value by writing a symbol, followed by an equals sign ‘=’, followed by an expression (see Chapter 6 [Expressions], page 29). This is equivalent to using the .set directive. See Section 7.43 [.set], page 43.

5.3 Symbol Names

Symbol names begin with a letter or with one of ‘._’. On most machines, you can also use $ in symbol names; exceptions are noted in Chapter 8 [Machine Dependencies], page 49. That character may be followed by any string of digits, letters, dollar signs (unless otherwise noted in Chapter 8 [Machine Dependencies], page 49), and underscores. For the AMD 29K family, ‘?’ is also allowed in the body of a symbol name, though not at its beginning.

Case of letters is significant: foo is a different symbol name than Foo.

Each symbol has exactly one name. Each name in an assembly language program refers to exactly one symbol. You may use that symbol name any number of times in a program.

Local Symbol Names

Local symbols help compilers and programmers use names temporarily. There are ten local symbol names, which are re-used throughout the program. You may refer to them using the names ‘0’ ‘1’ … ‘9’. To define a local symbol, write a label of the form ‘N:’ (where N represents any digit). To refer to the most recent previous definition of that symbol write ‘Nb’, using the same digit as when you defined the label. To refer to the next definition of a local label, write ‘Nf’—where N gives you a choice of 10 forward references. The ‘b’ stands for “backwards” and the ‘f’ stands for “forwards”.

Local symbols are not emitted by the current GNU C compiler.

There is no restriction on how you can use these labels, but remember that at any point in the assembly you can refer to at most 10 prior local labels and to at most 10 forward local labels.

Local symbol names are only a notation device. They are immediately transformed into more conventional symbol names before the assembler uses them. The symbol names stored in the symbol table, appearing in error messages and optionally emitted to the object file have these parts:

L All local labels begin with ‘L’. Normally both as and ld forget symbols that start with ‘L’. These labels are used for symbols you are never intended to see. If you give the ‘-L’ option then as will retain these symbols in the object file. If you also instruct ld to retain these symbols, you may use them in debugging.

digit If the label is written ‘0:’ then the digit is ‘0’. If the label is written ‘1:’ then the digit is ‘1’. And so on up through ‘9:’.

^A This unusual character is included so you don’t accidentally invent a symbol of the same name. The character has ASCII value ‘\001’.

ordinal number

This is a serial number to keep the labels distinct. The first ‘0:’ gets the number ‘1’; The 15th ‘0:’ gets the number ‘15’; etc.. Likewise for the other labels ‘1:’ through ‘9:’.

For instance, the first 1: is named L1^A1, the 44th 3: is named L3^A44.

5.4 The Special Dot Symbol

The special symbol ‘.’ refers to the current address that as is assembling into. Thus, the expression ‘melvin: .long .’ will cause melvin to contain its own address. Assigning a value to . is treated the same as a .org directive. Thus, the expression ‘.=.+4’ is the same as saying ‘.space 4’.

5.5 Symbol Attributes

Every symbol has, as well as its name, the attributes “Value” and “Type”. Depending on output format, symbols can also have auxiliary attributes.

If you use a symbol without defining it, as assumes zero for all these attributes, and probably won’t warn you. This makes the symbol an externally defined symbol, which is generally what you would want.

5.5.1 Value

The value of a symbol is (usually) 32 bits. For a symbol which labels a location in the text, data, bss or absolute sections the value is the number of addresses from the start of that section to the label. Naturally for text, data and bss sections the value of a symbol changes as ld changes section base addresses during linking. Absolute symbols’ values do not change during linking: that is why they are called absolute.

The value of an undefined symbol is treated in a special way. If it is 0 then the symbol is not defined in this assembler source program, and ld will try to determine its value from other programs it is linked with. You make this kind of symbol simply by mentioning a symbol name without defining it. A non-zero value represents a .comm common declaration. The value is how much common storage to reserve, in bytes (addresses). The symbol refers to the first address of the allocated storage.

5.5.2 Type

The type attribute of a symbol contains relocation (section) information, any flag settings indicating that a symbol is external, and (optionally), other information for linkers and debuggers. The exact format depends on the object-code output format in use.

28 Using as

5.5.3 Symbol Attributes: a.out 5.5.3.1 Descriptor

This is an arbitrary 16-bit value. You may establish a symbol’s descriptor value by using a .desc statement (see Section 7.11 [.desc], page 35). A descriptor value means nothing to as.

5.5.3.2 Other

This is an arbitrary 8-bit value. It means nothing to as.

5.5.4 Symbol Attributes for COFF

The COFF format supports a multitude of auxiliary symbol attributes; like the primary symbol attributes, they are set between .def and .endef directives.

5.5.4.1 Primary Attributes

The symbol name is set with .def; the value and type, respectively, with .val and .type.

5.5.4.2 Auxiliary Attributes

The as directives .dim, .line, .scl, .size, and .tag can generate auxiliary symbol table information for COFF.

6 Expressions

An expression specifies an address or numeric value. Whitespace may precede and/or follow an expression.

6.1 Empty Expressions

An empty expression has no value: it is just whitespace or null. Wherever an absolute expression is required, you may omit the expression and as will assume a value of (absolute) 0. This is compatible with other assemblers.

6.2 Integer Expressions

An integer expression is one or more arguments delimited by operators.

6.2.1 Arguments

Arguments are symbols, numbers or subexpressions. In other contexts arguments are sometimes called “arithmetic operands”. In this manual, to avoid confusing them with the “instruction operands” of the machine language, we use the term “argument” to refer to parts of expressions only, reserving the word “operand” to refer only to machine instruction operands.

Symbols are evaluated to yield {section NNN} where section is one of text, data, bss, absolute, or undefined. NNN is a signed, 2’s complement 32 bit integer.

Numbers are usually integers.

A number can be a flonum or bignum. In this case, you are warned that only the low order 32 bits are used, and as pretends these 32 bits are an integer. You may write integer-manipulating instructions that act on exotic constants, compatible with other assemblers.

Subexpressions are a left parenthesis ‘(’ followed by an integer expression, followed by a right parenthesis ‘)’; or a prefix operator followed by an argument.

6.2.2 Operators

Operators are arithmetic functions, like + or %. Prefix operators are followed by an argument. Infix operators appear between their arguments. Operators may be preceded and/or followed by whitespace.

6.2.3 Prefix Operator

as has the following prefix operators. They each take one argument, which must be absolute.

- Negation. Two’s complement negation. ~ Complementation. Bitwise not.

6.2.4 Infix Operators

Infix operators take two arguments, one on either side. Operators have precedence, but operations with equal precedence are performed left to right. Apart from + or -, both arguments must be absolute, and the result is absolute.

1. Highest Precedence * Multiplication. / Division. Truncation is the same as the C operator ‘/’ % Remainder.

<

<< Shift Left. Same as the C operator ‘<<’.

>

>> Shift Right. Same as the C operator ‘>>’.

2. Intermediate precedence

|

Bitwise Inclusive Or.

& Bitwise And. ^Bitwise Exclusive Or. ! Bitwise Or Not.

3. Lowest Precedence + Addition. If either argument is absolute, the result has the section of the other

argument. If either argument is pass1 or undefined, the result is pass1. Otherwise + is illegal.

- Subtraction. If the right argument is absolute, the result has the section of the left argument. If either argument is pass1 the result is pass1. If either argument is undefined the result is difference section. If both arguments are in the same section, the result is absolute—provided that section is one of text, data or bss. Otherwise subtraction is illegal.

The sense of the rule for addition is that it’s only meaningful to add the offsets in an address; you can only have a defined section in one of the two arguments.

Similarly, you can’t subtract quantities from two different sections.

32 Using as

7 Assembler Directives

All assembler directives have names that begin with a period (‘.’). The rest of the name is letters, usually in lower case.

This chapter discusses directives that are available regardless of the target machine configuration for the GNU assembler. Some machine configurations provide additional directives. See Chapter 8 [Machine Dependencies], page 49.

7.1 .abort

This directive stops the assembly immediately. It is for compatibility with other assemblers. The original idea was that the assembly language source would be piped into the assembler. If the sender of the source quit, it could use this directive tells as to quit also. One day .abort will not be supported.

7.2 .ABORT

When producing COFF output, as accepts this directive as a synonym for ‘.abort’.

When producing b.out output, as accepts this directive, but ignores it.

7.3 .align abs-expr , abs-expr

Pad the location counter (in the current subsection) to a particular storage boundary. The first expression (which must be absolute) is the number of low-order zero bits the location counter will have after advancement. For example ‘.align 3’ will advance the location counter until it a multiple of 8. If the location counter is already a multiple of 8, no change is needed.

The second expression (also absolute) gives the value to be stored in the padding bytes. It (and the comma) may be omitted. If it is omitted, the padding bytes are zero.

7.4 .app-file string

.app-file (which may also be spelled ‘.file’) tells as that we are about to start a new logical file. string is the new file name. In general, the filename is recognized whether or not it is surrounded by quotes ‘"’; but if you wish to specify an empty file name is permitted, you must give the quotes–"". This statement may go away in future: it is only recognized to be compatible with old as programs.

7.5 .ascii "string"…

.ascii expects zero or more string literals (see Section 3.6.1.1 [Strings], page 14) separated by commas. It assembles each string (with no automatic trailing zero byte) into consecutive addresses.

7.6 .asciz "string"…

.asciz is just like .ascii, but each string is followed by a zero byte. The “z” in ‘.asciz’ stands for “zero”.

7.7 .byte expressions

.byte expects zero or more expressions, separated by commas. Each expression is assembled into the next byte.

7.8 .comm symbol , length

.comm declares a named common area in the bss section. Normally ld reserves memory addresses for it during linking, so no partial program defines the location of the symbol. Use .comm to tell ld that it must be at least length bytes long. ld will allocate space for each .comm symbol that is at least as long as the longest .comm request in any of the partial programs linked. length is an absolute expression.

7.9 .data subsection

.data tells as to assemble the following statements onto the end of the data subsection numbered subsection (which is an absolute expression). If subsection is omitted, it defaults to zero.

7.10 .def name

Begin defining debugging information for a symbol name; the definition extends until the .endef directive is encountered.

This directive is only observed when as is configured for COFF format output; when producing b.out, ‘.def’ is recognized, but ignored.

7.11 .desc symbol, abs-expression

This directive sets the descriptor of the symbol (see Section 5.5 [Symbol Attributes], page 27) to the low 16 bits of an absolute expression.

The ‘.desc’ directive is not available when as is configured for COFF output; it is only for a.out or b.out object format. For the sake of compatibility, as will accept it, but produce no output, when configured for COFF.

7.12 .dim

This directive is generated by compilers to include auxiliary debugging information in the symbol table. It is only permitted inside .def/.endef pairs.

‘.dim’ is only meaningful when generating COFF format output; when as is generating b.out, it accepts this directive but ignores it.

7.13 .double flonums

.double expects zero or more flonums, separated by commas. It assembles floating point numbers. The exact kind of floating point numbers emitted depends on how as is configured. See Chapter 8 [Machine Dependencies], page 49.

7.14 .eject

Force a page break at this point, when generating assembly listings.

7.15 .else

.else is part of the as support for conditional assembly; see Section 7.26 [.if], page 38. It marks the beginning of a section of code to be assembled if the condition for the preceding .if was false.

7.16 .endef

This directive flags the end of a symbol definition begun with .def.

‘.endef’ is only meaningful when generating COFF format output; if as is configured to generate b.out, it accepts this directive but ignores it.

7.17 .endif

.endif is part of the as support for conditional assembly; it marks the end of a block of code that is only assembled conditionally. See Section 7.26 [.if], page 38.

7.18 .equ symbol, expression

This directive sets the value of symbol to expression. It is synonymous with ‘.set’; see Section 7.43 [.set], page 43.

7.19 .extern

.extern is accepted in the source program—for compatibility with other assemblers—but it is ignored. as treats all undefined symbols as external.

7.20 .file string

.file (which may also be spelled ‘.app-file’) tells as that we are about to start a new logical file. string is the new file name. In general, the filename is recognized whether or not it is surrounded by quotes ‘"’; but if you wish to specify an empty file name, you must give the quotes–"". This statement may go away in future: it is only recognized to be compatible with old as programs. In some configurations of as, .file has already been removed to avoid conflicts with other assemblers. See Chapter 8 [Machine Dependencies], page 49.

7.21 .fill repeat , size , value

result, size and value are absolute expressions. This emits repeat copies of size bytes. Repeat may be zero or more. Size may be zero or more, but if it is more than 8, then it is deemed to have the value 8, compatible with other people’s assemblers. The contents of each repeat bytes is taken from an 8-byte number. The highest order 4 bytes are zero. The lowest order 4 bytes are value rendered in the byte-order of an integer on the computer as is assembling for. Each size bytes in a repetition is taken from the lowest order size bytes of this number. Again, this bizarre behavior is compatible with other people’s assemblers.

size and value are optional. If the second comma and value are absent, value is assumed zero. If the first comma and following tokens are absent, size is assumed to be 1.

7.22 .float flonums

This directive assembles zero or more flonums, separated by commas. It has the same effect as .single. The exact kind of floating point numbers emitted depends on how as is configured. See Chapter 8 [Machine Dependencies], page 49.

7.23 .global symbol, .globl symbol

.global makes the symbol visible to ld. If you define symbol in your partial program, its value is made available to other partial programs that are linked with it. Otherwise, symbol will take its attributes from a symbol of the same name from another partial program it is linked with.

Both spellings (‘.globl’ and ‘.global’) are accepted, for compatibility with other assemblers.

7.24 .hword expressions

This expects zero or more expressions, and emits a 16 bit number for each.

This directive is a synonym for ‘.short’; depending on the target architecture, it may also be a synonym for ‘.word’.

7.25 .ident

This directive is used by some assemblers to place tags in object files. as simply accepts the directive for source-file compatibility with such assemblers, but does not actually emit anything for it.

7.26 .if absolute expression

.if marks the beginning of a section of code which is only considered part of the source program being assembled if the argument (which must be an absolute expression) is non-zero. The end of the conditional section of code must be marked by .endif (see Section 7.17 [.endif], page 36); optionally, you may include code for the alternative condition, flagged by .else (see Section 7.15 [.else], page 36.

The following variants of .if are also supported:

.ifdef symbol

Assembles the following section of code if the specified symbol has been defined.

.ifndef symbol
ifnotdef symbol

Assembles the following section of code if the specified symbol has not been defined. Both spelling variants are equivalent.

7.27 .include "file"

This directive provides a way to include supporting files at specified points in your source program. The code from file is assembled as if it followed the point of the .include; when the end of the included file is reached, assembly of the original file continues. You can control the search paths used with the ‘-I’ command-line option (see Chapter 2 [Command-Line Options], page 7). Quotation marks are required around file.

7.28 .int expressions

Expect zero or more expressions, of any section, separated by commas. For each expression, emit a 32-bit number that will, at run time, be the value of that expression. The byte order of the expression depends on what kind of computer will run the program.

7.29 .lcomm symbol , length

Reserve length (an absolute expression) bytes for a local common denoted by symbol. The section and value of symbol are those of the new local common. The addresses are allocated in the bss section, so at run-time the bytes will start off zeroed. Symbol is not declared global (see Section 7.23 [.global], page 38), so is normally not visible to ld.

7.30 .lflags

as accepts this directive, for compatibility with other assemblers, but ignores it.

7.31 .line line-number

Tell as to change the logical line number. line-number must be an absolute expression. The next line will have that logical line number. So any other statements on the current line (after a statement separator character) will be reported as on logical line number line-number − 1. One day this directive will be unsupported: it is used only for compatibility with existing assembler programs.

Warning: In the AMD29K configuration of as, this command is only available with the name .ln, rather than as either .line or .ln.

Even though this is a directive associated with the a.out or b.out object-code formats, as will still recognize it when producing COFF output, and will treat ‘.line’ as though it were the COFF ‘.ln’ if it is found outside a .def/.endef pair.

Inside a .def, ‘.line’ is, instead, one of the directives used by compilers to generate auxiliary symbol information for debugging.

7.32 .ln line-number

‘.ln’ is a synonym for ‘.line’.

7.33 .list

Control (in conjunction with the .nolist directive) whether or not assembly listings are generated. These two directives maintain an internal counter (which is zero initially). .list increments the counter, and .nolist decrements it. Assembly listings are generated whenever the counter is greater than zero.

By default, listings are disabled. When you enable them (with the ‘-a’ command line option; see Chapter 2 [Command-Line Options], page 7), the initial value of the listing counter is one.

7.34 .long expressions

.long is the same as ‘.int’, see Section 7.28 [.int], page 39.

7.35 .nolist

Control (in conjunction with the .list directive) whether or not assembly listings are generated. These two directives maintain an internal counter (which is zero initially). .list increments the counter, and .nolist decrements it. Assembly listings are generated whenever the counter is greater than zero.

7.36 .octa bignums

This directive expects zero or more bignums, separated by commas. For each bignum, it emits a 16-byte integer.

The term “octa” comes from contexts in which a “word” is two bytes; hence octa-word for 16 bytes.

7.37 .org new-lc , fill

.org will advance the location counter of the current section to new-lc. new-lc is either an absolute expression or an expression with the same section as the current subsection. That is, you can’t use .org to cross sections: if new-lc has the wrong section, the .org directive is ignored. To be compatible with former assemblers, if the section of new-lc is absolute, as will issue a warning, then pretend the section of new-lc is the same as the current subsection.

.org may only increase the location counter, or leave it unchanged; you cannot use .org to move the location counter backwards.

Because as tries to assemble programs in one pass new-lc may not be undefined. If you really detest this restriction we eagerly await a chance to share your improved assembler.

Beware that the origin is relative to the start of the section, not to the start of the subsection. This is compatible with other people’s assemblers.

When the location counter (of the current subsection) is advanced, the intervening bytes are filled with fill which should be an absolute expression. If the comma and fill are omitted, fill defaults to zero.

7.38 .psize lines , columns

Use this directive to declare the number of lines—and, optionally, the number of columns—to use for each page, when generating listings.

If you don’t use .psize, listings will use a default line-count of 60. You may omit the comma and columns specification; the default width is 200 columns.

as will generate formfeeds whenever the specified number of lines is exceeded (or whenever you explicitly request one, using .eject).

If you specify lines as 0, no formfeeds are generated save those explicitly specified with .eject.

7.39 .quad bignums

.quad expects zero or more bignums, separated by commas. For each bignum, it emits an 8-byte integer. If the bignum won’t fit in 8 bytes, it prints a warning message; and just takes the lowest order 8 bytes of the bignum.

The term “quad” comes from contexts in which a “word” is two bytes; hence quad-word for 8 bytes.

7.40 .sbttl "subheading"

Use subheading as the title (third line, immediately after the title line) when generating assembly listings.

This directive affects subsequent pages, as well as the current page if it appears within ten lines of the top of a page.

7.41 .scl class

Set the storage-class value for a symbol. This directive may only be used inside a .def/.endef pair. Storage class may flag whether a symbol is static or external, or it may record further symbolic debugging information.

The ‘.scl’ directive is primarily associated with COFF output; when configured to generate b.out output format, as will accept this directive but ignore it.

7.42 .section name, subsection

Assemble the following code into end of subsection numbered subsection in the COFF named section name. If you omit subsection, as uses subsection number zero. ‘.section .text’ is equivalent to the .text directive; ‘.section .data’ is equivalent to the .data directive.

7.43 .set symbol, expression

This directive sets the value of symbol to expression. This will change symbol’s value and type to conform to expression. If symbol was flagged as external, it remains flagged. (See Section 5.5 [Symbol Attributes], page 27.)

You may .set a symbol many times in the same assembly. If the expression’s section is unknowable during pass 1, a second pass over the source program will be forced. The second pass is currently not implemented. as will abort with an error message if one is required.

If you .set a global symbol, the value stored in the object file is the last value stored into it.

7.44 .short expressions

.short is normally the same as ‘.word’. See Section 7.54 [.word], page 46.

In some configurations, however, .short and .word generate numbers of different lengths; see Chapter 8 [Machine Dependencies], page 49.

7.45 .single flonums

This directive assembles zero or more flonums, separated by commas. It has the same effect as .float. The exact kind of floating point numbers emitted depends on how as is configured. See Chapter 8 [Machine Dependencies], page 49.

7.46 .size

This directive is generated by compilers to include auxiliary debugging information in the symbol table. It is only permitted inside .def/.endef pairs.

‘.size’ is only meaningful when generating COFF format output; when as is generating b.out, it accepts this directive but ignores it.

7.47 .space size , fill

This directive emits size bytes, each of value fill. Both size and fill are absolute expressions. If the comma and fill are omitted, fill is assumed to be zero.

On the AMD 29K, this directive is ignored; it is accepted for compatibility with other AMD 29K assemblers.

Warning: In most versions of the GNU assembler, the directive .space has the effect of .block See Chapter 8 [Machine Dependencies], page 49.

7.48 .stabd, .stabn, .stabs

There are three directives that begin ‘.stab’. All emit symbols (see Chapter 5 [Symbols], page 25), for use by symbolic debuggers. The symbols are not entered in the as hash table: they cannot be referenced elsewhere in the source file. Up to five fields are required:

string This is the symbol’s name. It may contain any character except ‘\000’, so is more general than ordinary symbol names. Some debuggers used to code arbitrarily complex structures into symbol names using this field.

type An absolute expression. The symbol’s type is set to the low 8 bits of this expression.

Any bit pattern is permitted, but ld and debuggers will choke on silly bit patterns.

other An absolute expression. The symbol’s “other” attribute is set to the low 8 bits of this

expression.

desc An absolute expression. The symbol’s descriptor is set to the low 16 bits of this

expression.

value An absolute expression which becomes the symbol’s value.

If a warning is detected while reading a .stabd, .stabn, or .stabs statement, the symbol has probably already been created and you will get a half-formed symbol in your object file. This is compatible with earlier assemblers!

.stabd type , other , desc The “name” of the symbol generated is not even an empty string. It is a null pointer, for compatibility. Older assemblers used a null pointer so they didn’t waste space in object files with empty strings. The symbol’s value is set to the location counter, relocatably. When your program is linked, the value of this symbol will be where the location counter was when the .stabd was assembled.

.stabn type , other , desc , value

The name of the symbol is set to the empty string "".

.stabs string , type , other , desc , value All five fields are specified.

7.49 .tag structname

This directive is generated by compilers to include auxiliary debugging information in the symbol table. It is only permitted inside .def/.endef pairs. Tags are used to link structure definitions in the symbol table with instances of those structures.

‘.tag’ is only used when generating COFF format output; when as is generating b.out, it accepts this directive but ignores it.

7.50 .text subsection

Tells as to assemble the following statements onto the end of the text subsection numbered subsection, which is an absolute expression. If subsection is omitted, subsection number zero is used.

7.51 .title "heading"

Use heading as the title (second line, immediately after the source file name and pagenumber) when generating assembly listings.

This directive affects subsequent pages, as well as the current page if it appears within ten lines of the top of a page.

7.52 .type int

This directive, permitted only within .def/.endef pairs, records the integer int as the type attribute of a symbol table entry.

‘.type’ is associated only with COFF format output; when as is configured for b.out output, it accepts this directive but ignores it.

7.53 .val addr

This directive, permitted only within .def/.endef pairs, records the address addr as the value attribute of a symbol table entry.

‘.val’ is used only for COFF output; when as is configured for b.out, it accepts this directive but ignores it.

7.54 .word expressions

This directive expects zero or more expressions, of any section, separated by commas.

The size of the number emitted, and its byte order, depends on what kind of computer will run the program.

Warning: Special Treatment to support Compilers

Machines with a 32-bit address space, but that do less than 32-bit addressing, require the following special treatment. If the machine of interest to you does 32-bit addressing (or doesn’t require it; see Chapter 8 [Machine Dependencies], page 49), you can ignore this issue.

In order to assemble compiler output into something that will work, as will occasionlly do strange things to ‘.word’ directives. Directives of the form ‘.word sym1-sym2’ are often emitted by compilers as part of jump tables. Therefore, when as assembles a directive of the form ‘.word sym1-sym2’, and the difference between sym1 and sym2 does not fit in 16 bits, as will create a secondary jump table, immediately before the next label. This secondary jump table will be preceded by a short-jump to the first byte after the secondary table. This short-jump prevents the flow of control from accidentally falling into the new table. Inside the table will be a long-jump to sym2. The original ‘.word’ will contain sym1 minus the address of the long-jump to sym2.

If there were several occurrences of ‘.word sym1-sym2’ before the secondary jump table, all of them will be adjusted. If there was a ‘.word sym3-sym4’, that also did not fit in sixteen bits, a long-jump to sym4 will be included in the secondary jump table, and the .word directives will be adjusted to contain sym3 minus the address of the long-jump to sym4; and so on, for as many entries in the original jump table as necessary.

7.55 Deprecated Directives

One day these directives won’t work. They are included for compatibility with older assemblers.

.abort
.app-file
.line

48 Using as

8 Machine Dependent Features

The machine instruction sets are (almost by definition) different on each machine where as runs. Floating point representations vary as well, and as often supports a few additional directives or command-line options for compatibility with other assemblers on a particular platform. Finally, some versions of as support special pseudo-instructions for branch optimization.

This chapter discusses most of these differences, though it does not include details on any machine’s instruction set. For details on that subject, see the hardware manufacturer’s manual.

8.1 VAX Dependent Features

8.1.1 VAX Command-Line Options

The Vax version of as accepts any of the following options, gives a warning message that the option was ignored and proceeds. These options are for compatibility with scripts designed for other people’s assemblers.

-D (Debug) -S (Symbol Table) -T (Token Trace)

These are obsolete options used to debug old assemblers.

-d (Displacement size for JUMPs)

This option expects a number following the -d. Like options that expect filenames, the number may immediately follow the -d (old standard) or constitute the whole of the command line argument that follows -d (GNU standard).

-V (Virtualize Interpass Temporary File)

Some other assemblers use a temporary file. This option commanded them to keep the information in active memory rather than in a disk file. as always does this, so this option is redundant.

-J (JUMPify Longer Branches)

Many 32-bit computers permit a variety of branch instructions to do the same job. Some of these instructions are short (and fast) but have a limited range; others are long (and slow) but can branch anywhere in virtual memory. Often there are 3 flavors of branch: short, medium and long. Some other assemblers would emit short and medium branches, unless told by this option to emit short and long branches.

-t (Temporary File Directory)

Some other assemblers may use a temporary file, and this option takes a filename being the directory to site the temporary file. Since as does not use a temporary disk file, this option makes no difference. -t needs exactly one filename.

The Vax version of the assembler accepts two options when compiled for VMS. They are -h, and -+. The -h option prevents as from modifying the symbol-table entries for symbols that contain lowercase characters (I think). The -+ option causes as to print warning messages if the FILENAME part of the object file, or any symbol name is larger than 31 characters. The -+ option also insertes some code following the ‘_main’ symbol so that the object file will be compatible with Vax-11 "C".

8.1.2 VAX Floating Point

Conversion of flonums to floating point is correct, and compatible with previous assemblers. Rounding is towards zero if the remainder is exactly half the least significant bit.

D, F, G and H floating point formats are understood.

Immediate floating literals (e.g. ‘S`$6.9’) are rendered correctly. Again, rounding is towards zero in the boundary case.

The .float directive produces f format numbers. The .double directive produces d format numbers.

8.1.3 Vax Machine Directives

The Vax version of the assembler supports four directives for generating Vax floating point constants. They are described in the table below.

.dfloat This expects zero or more flonums, separated by commas, and assembles Vax d format

64-bit floating point constants.

.ffloat This expects zero or more flonums, separated by commas, and assembles Vax f format 32-bit floating point constants.

.gfloat This expects zero or more flonums, separated by commas, and assembles Vax g format

64-bit floating point constants.

.hfloat This expects zero or more flonums, separated by commas, and assembles Vax h format 128-bit floating point constants.

8.1.4 VAX Opcodes

All DEC mnemonics are supported. Beware that case… instructions have exactly 3 operands. The dispatch table that follows the case… instruction should be made with .word statements. This is compatible with all unix assemblers we know of.

8.1.5 VAX Branch Improvement

Certain pseudo opcodes are permitted. They are for branch instructions. They expand to the shortest branch instruction that will reach the target. Generally these mnemonics are made by substituting ‘j’ for ‘b’ at the start of a DEC mnemonic. This feature is included both for compatibility and to help compilers. If you don’t need this feature, don’t use these opcodes. Here are the mnemonics, and the code they can expand into.

jbsb ‘Jsb’ is already an instruction mnemonic, so we chose ‘jbsb’.

(byte displacement)

                    bsbb . . .

(word displacement)

                    bsbw . . .

(long displacement)

                    jsb . . .
jbr

jr Unconditional branch.

(byte displacement)

                    brb . . .

(word displacement)

                    brw . . .

(long displacement)

                    jmp . . .

jCOND COND may be any one of the conditional branches neq, nequ, eql, eqlu, gtr, geq, lss, gtru, lequ, vc, vs, gequ, cc, lssu, cs. COND may also be one of the bit tests bs, bc, bss, bcs, bsc, bcc, bssi, bcci, lbs, lbc. NOTCOND is the opposite condition to COND. (byte displacement)

bCOND …

(word displacement)

          bNOTCOND foo ; brw . . . ; foo:

(long displacement)

          bNOTCOND foo ; jmp . . . ; foo:

jacbX X may be one of b d f g h l w.

(word displacement)

OPCODE …

(long displacement)

OPCODE …, foo ;

               brb bar ;
               foo: jmp . . . ;
               bar:

jaobYYY YYY may be one of lss leq. jsobZZZ ZZZ may be one of geq gtr.

(byte displacement)

OPCODE …

(word displacement)

OPCODE …, foo ;

               brb bar ;

foo: brw destination ;

               bar:

(long displacement)

OPCODE …, foo ;

               brb bar ;

foo: jmp destination ;

               bar:
aobleq
aoblss
sobgeq
sobgtr

(byte displacement)

OPCODE … (word displacement)

OPCODE …, foo ;

brb bar ;

foo: brw destination ;

bar:

(long displacement) OPCODE …, foo ;

brb bar ;

foo: jmp destination ;

bar:

8.1.6 VAX Operands

The immediate character is ‘$’ for Unix compatibility, not ‘#’ as DEC writes it. The indirect character is ‘*’ for Unix compatibility, not ‘@’ as DEC writes it. The displacement sizing character is ‘`’ (an accent grave) for Unix compatibility, not ‘^’ as DEC

writes it. The letter preceding ‘`’ may have either case. ‘G’ is not understood, but all other letters (b i l s w) are understood.

Register names understood are r0 r1 r2 … r15 ap fp sp pc. Any case of letters will do. For instance

tstb *w`$4(r5)

Any expression is permitted in an operand. Operands are comma separated.

8.1.7 Not Supported on VAX

Vax bit fields can not be assembled with as. Someone can add the required code if they really need it.

8.2 AMD 29K Dependent Features

8.2.1 Options

as has no additional command-line options for the AMD 29K family.

8.2.2 Syntax

8.2.2.1 Special Characters

‘;’ is the line comment character.

‘@’ can be used instead of a newline to separate statements.

The character ‘?’ is permitted in identifiers (but may not begin an identifier).

8.2.2.2 Register Names

General-purpose registers are represented by predefined symbols of the form ‘GRnnn’ (for global registers) or ‘LRnnn’ (for local registers), where nnn represents a number between 0 and 127, written with no leading zeros. The leading letters may be in either upper or lower case; for example, ‘gr13’ and ‘LR7’ are both valid register names.

You may also refer to general-purpose registers by specifying the register number as the result of an expression (prefixed with ‘%%’ to flag the expression as a register number):

%%expression

—where expression must be an absolute expression evaluating to a number between 0 and 255. The range [0, 127] refers to global registers, and the range [128, 255] to local registers.

In addition, as understands the following protected special-purpose register names for the AMD 29K family:

vab    chd    pc0
ops    chc    pc1
cps    rbp    pc2
cfg    tmc    mmu
cha    tmr    lru

These unprotected special-purpose register names are also recognized:

ipc    alu    fpe
ipa    bp     inte
ipb    fc     fps
q      cr     exop

8.2.3 Floating Point

The AMD 29K family uses ieee floating-point numbers.

8.2.4 AMD 29K Machine Directives

.block size , fill

This directive emits size bytes, each of value fill. Both size and fill are absolute expressions. If the comma and fill are omitted, fill is assumed to be zero.

In other versions of the GNU assembler, this directive is called ‘.space’.

.cputype This directive is ignored; it is accepted for compatibility with other AMD 29K assemblers.

.file This directive is ignored; it is accepted for compatibility with other AMD 29K assemblers.

Warning: in other versions of the GNU assembler, .file is used for the directive called .app-file in the AMD 29K support.

.line This directive is ignored; it is accepted for compatibility with other AMD 29K assemblers.

.sect This directive is ignored; it is accepted for compatibility with other AMD 29K assemblers.

.use section name

Establishes the section and subsection for the following code; section name may be one of .text, .data, .data1, or .lit. With one of the first three section name options, ‘.use’ is equivalent to the machine directive section name; the remaining case, ‘.use .lit’, is the same as ‘.data 200’.

8.2.5 Opcodes

as implements all the standard AMD 29K opcodes. No additional pseudo-instructions are needed on this family.

For information on the 29K machine instruction set, see Am29000 User’s Manual, Advanced Micro Devices, Inc.

8.3 H8/300 Dependent Features

8.3.1 Options

as has no additional command-line options for the Hitachi H8/300 family.

8.3.2 Syntax

8.3.2.1 Special Characters

‘;’ is the line comment character.

‘$’ can be used instead of a newline to separate statements. Therefore you may not use ‘$’ in symbol names on the H8/300.

8.3.2.2 Register Names

You can use predefined symbols of the form ‘rnh’ and ‘rnl’ to refer to the H8/300 registers as sixteen 8-bit general-purpose registers. n is a digit from ‘0’ to ‘7’); for instance, both ‘r0h’ and ‘r7l’ are valid register names.

You can also use the eight predefined symbols ‘rn’ to refer to the H8/300 registers as 16-bit registers (you must use this form for addressing).

The two control registers are called pc (program counter; a 16-bit register) and ccr (condition code register; an 8-bit register). r7 is used as the stack pointer, and can also be called sp.

8.3.2.3 Addressing Modes

as understands the following addressing modes for the H8/300: rn Register direct @rn Register indirect

@(d, rn)
@(d:16, rn)

Register indirect: 16-bit displacement d from register n. (You may specify the ‘:16’ for clarity if you wish, but it is not required and has no effect.)

@rn+ Register indirect with post-increment @-rn Register indirect with pre-decrement @aa

@aa:8

@aa:16 Absolute address aa. You may specify the ‘:8’ or ‘:16’ for clarity, if you wish; but as

neither requires this nor uses it—the address size required is taken from context.

#xx

#xx:8

#xx:16 Immediate data xx. You may specify the ‘:8’ or ‘:16’ for clarity, if you wish; but as

neither requires this nor uses it—the data size required is taken from context.

@@aa

@@aa:8 Memory indirect. You may specify the ‘:8’ for clarity, if you wish; but as neither requires this nor uses it.

8.3.3 Floating Point

The H8/300 family uses ieee floating-point numbers.

8.3.4 H8/300 Machine Directives

as has no machine-dependent directives for the H8/300. However, on this platform the ‘.int’ and ‘.word’ directives generate 16-bit numbers.

8.3.5 Opcodes

For detailed information on the H8/300 machine instruction set, see H8/300 Series Programming Manual (Hitachi ADE–602–025).

as implements all the standard H8/300 opcodes. No additional pseudo-instructions are needed on this family.

Chapter 8: Machine Dependent Features 59 The following table summarizes the opcodes and their arguments:

Legend: Rs source register Rd destination register imm immediate data x:3 a bit (as a number between 0 and 7) d:8 eight bit displacement from pc d:16 sixteen bit displacement from Rs

add.b   Rs,Rd                  biand   #x:3,Rd
add.b   #imm:8,Rd              biand   #x:3,@Rd
add.w   Rs,Rd                  biand   #x:3,@aa:8
adds    #1,Rd                  bild    #x:3,Rd
adds    #2,Rd                  bild    #x:3,@Rd
addx    #imm:8,Rd              bild    #x:3,@aa:8
addx    Rs,Rd                  bior    #x:3,Rd
and     #imm:8,Rd              bior    #x:3,@Rd
and     Rs,Rd                  bior    #x:3,@aa:8
andc    #imm:8,ccr             bist    #x:3,Rd
band    #x:3,Rd                bist    #x:3,@Rd
band    #x:3,@Rd               bist    #x:3,@aa:8
bra     d:8                    bixor   #x:3,Rd
bt      d:8                    bixor   #x:3,@Rd
brn     d:8                    bixor   #x:3,@aa:8
bf      d:8                    bld     #x:3,Rd
bhi     d:8                    bld     #x:3,@Rd
bls     d:8                    bld     #x:3,@aa:8
bcc     d:8                    bnot    #x:3,Rd
bhs     d:8                    bnot    #x:3,@Rd
bcs     d:8                    bnot    #x:3,@aa:8
blo     d:8                    bnot    Rs,Rd
bne     d:8                    bnot    Rs,@Rd
beq     d:8                    bnot    Rs,@aa:8
bvc     d:8                    bor     #x:3,Rd
bvs     d:8                    bor     #x:3,@Rd
bpl     d:8                    bor     #x:3,@aa:8
bmi     d:8                    bset    #x:3,@Rd
bge     d:8                    bset    #x:3,@aa:8
blt     d:8                    bset    Rs,Rd
bgt     d:8                    bset    Rs,@Rd
ble     d:8                    bset    Rs,@aa:8
bclr    #x:3,Rd                bsr     d:8
bclr    #x:3,@Rd               bst     #x:3,Rd
bclr    #x:3,@aa:8             bst     #x:3,@Rd
bclr    Rs,Rd                  bst     #x:3,@aa:8
bclr    Rs,@Rd                 btst    #x:3,Rd

btst    #x:3,@Rd               mov.w   @(d:16, Rs),Rd
btst    #x:3,@aa:8             mov.w   @Rs+,Rd
btst    Rs,Rd                  mov.w   @aa:16,Rd
btst    Rs,@Rd                 mov.w   Rs,@Rd
btst    Rs,@aa:8               mov.w   Rs,@(d:16, Rd)
bxor    #x:3,Rd                mov.w   Rs,@-Rd
bxor    #x:3,@Rd               mov.w   Rs,@aa:16
bxor    #x:3,@aa:8             movfpe  @aa:16,Rd
cmp.b   #imm:8,Rd              movtpe  Rs,@aa:16
cmp.b   Rs,Rd                  mulxu   Rs,Rd
cmp.w   Rs,Rd                  neg     Rs
daa     Rs                     nop
das     Rs                     not     Rs
dec     Rs                     or      #imm:8,Rd
divxu   Rs,Rd                  or      Rs,Rd
eepmov                         orc     #imm:8,ccr
inc     Rs                     pop     Rs
jmp     @Rs                    push    Rs
jmp     @aa:16                 rotl    Rs
jmp     @@aa                   rotr    Rs
jsr     @Rs                    rotxl   Rs
jsr     @aa:16                 rotxr   Rs
jsr     @@aa:8                 rte
ldc     #imm:8,ccr             rts
ldc     Rs,ccr                 shal    Rs
mov.b   Rs,Rd                  shar    Rs
mov.b   #imm:8,Rd              shll    Rs
mov.b   @Rs,Rd                 shlr    Rs
mov.b   @(d:16, Rs),Rd         sleep
mov.b   @Rs+,Rd                stc     ccr,Rd
mov.b   @aa:16,Rd              sub.b   Rs,Rd
mov.b   @aa:8,Rd               sub.w   Rs,Rd
mov.b   Rs,@Rd                 subs    #1,Rd
mov.b   Rs,@(d:16, Rd)         subs    #2,Rd
mov.b   Rs,@-Rd                subx    #imm:8,Rd
mov.b   Rs,@aa:16              subx    Rs,Rd
mov.b   Rs,@aa:8               xor     #imm:8,Rd
mov.w   Rs,Rd                  xor     Rs,Rd
mov.w   #imm:16,Rd             xorc    #imm:8,ccr
mov.w   @Rs,Rd

Four H8/300 instructions (add, cmp, mov, sub) are defined with variants using the suffixes ‘.b’ and ‘.w’ to specify the size of a memory operand. as supports these suffixes, but does not require them; since one of the operands is always a register, as can deduce the correct size.

For example, since r0 refers to a 16-bit register,

     mov    r0,@foo

is equivalent to

     mov.w  r0,@foo

If you use the size suffixes, as will issue a warning if there’s a mismatch between the suffix and the register size.

8.4 H8/500 Dependent Features

8.4.1 Options

as has no additional command-line options for the Hitachi H8/500 family.

8.4.2 Syntax

8.4.2.1 Special Characters

‘!’ is the line comment character.

‘;’ can be used instead of a newline to separate statements.

Since ‘$’ has no special meaning, you may use it in symbol names.

8.4.2.2 Register Names

You can use the predefined symbols ‘r0’, ‘r1’, ‘r2’, ‘r3’, ‘r4’, ‘r5’, ‘r6’, and ‘r7’ to refer to the H8/500 registers.

The H8/500 also has these control registers: cp code pointer dp data pointer bp base pointer tp stack top pointer ep extra pointer sr status register ccr condition code register

All registers are 16 bits long. To represent 32 bit numbers, use two adjacent registers; for distant memory addresses, use one of the segment pointers (cp for the program counter; dp for r0–r3; ep for r4 and r5; and tp for r6 and r7.

8.4.2.3 Addressing Modes

as understands the following addressing modes for the H8/500:

Rn Register direct @Rn Register indirect

@(d:8, Rn)

Register indirect with 8 bit signed displacement

@(d:16, Rn)

Register indirect with 16 bit signed displacement

@-Rn Register indirect with pre-decrement @Rn+ Register indirect with post-increment @aa:8 8 bit absolute address @aa:16 16 bit absolute address #xx:8 8 bit immediate #xx:16 16 bit immediate

8.4.3 Floating Point

The H8/500 family uses ieee floating-point numbers.

8.4.4 H8/500 Machine Directives

as has no machine-dependent directives for the H8/500. However, on this platform the ‘.int’ and ‘.word’ directives generate 16-bit numbers.

8.4.5 Opcodes

For detailed information on the H8/500 machine instruction set, see H8/500 Series Programming Manual (Hitachi M21T001).

as implements all the standard H8/500 opcodes. No additional pseudo-instructions are needed on this family.

The following table summarizes H8/500 opcodes and their operands:

Legend: abs8 8-bit absolute address abs16 16-bit absolute address abs24 24-bit absolute address

crb       ccr, br, ep, dp, tp, dp

disp8 8-bit displacement

ea        rn, @rn, @(d:8, rn), @(d:16, rn),
          @-rn, @rn+, @aa:8, @aa:16,
          #xx:8, #xx:16
ea_mem    @rn, @(d:8, rn), @(d:16, rn),
          @-rn, @rn+, @aa:8, @aa:16
ea_noimm  rn, @rn, @(d:8, rn), @(d:16, rn),
          @-rn, @rn+, @aa:8, @aa:16
fp        r6

imm4 4-bit immediate data imm8 8-bit immediate data imm16 16-bit immediate data pcrel8 8-bit offset from program counter pcrel16 16-bit offset from program counter

qim       -2, -1, 1, 2

rd any register rs a register distinct from rd rlist comma-separated list of registers in parentheses;

register ranges rd-rs are allowed

sp stack pointer (r7) sr status register sz size; ‘.b’ or ‘.w’. If omitted, default ‘.w’

ldc[.b] ea,crb                 bcc[.w] pcrel16
ldc[.w] ea,sr                  bcc[.b] pcrel8
add[:q] sz qim,ea_noimm        bhs[.w] pcrel16
add[:g] sz ea,rd               bhs[.b] pcrel8
adds sz ea,rd                  bcs[.w] pcrel16
addx sz ea,rd                  bcs[.b] pcrel8
and sz ea,rd                   blo[.w] pcrel16
andc[.b] imm8,crb              blo[.b] pcrel8
andc[.w] imm16,sr              bne[.w] pcrel16
bpt                            bne[.b] pcrel8
bra[.w] pcrel16                beq[.w] pcrel16
bra[.b] pcrel8                 beq[.b] pcrel8
bt[.w] pcrel16                 bvc[.w] pcrel16
bt[.b] pcrel8                  bvc[.b] pcrel8
brn[.w] pcrel16                bvs[.w] pcrel16
brn[.b] pcrel8                 bvs[.b] pcrel8
bf[.w] pcrel16                 bpl[.w] pcrel16
bf[.b] pcrel8                  bpl[.b] pcrel8
bhi[.w] pcrel16                bmi[.w] pcrel16
bhi[.b] pcrel8                 bmi[.b] pcrel8
bls[.w] pcrel16                bge[.w] pcrel16
bls[.b] pcrel8                 bge[.b] pcrel8

Chapter 8: Machine Dependent Features 67

blt[.w] pcrel16                mov[:g][.b] imm8,ea_mem
blt[.b] pcrel8                 mov[:g][.w] imm16,ea_mem
bgt[.w] pcrel16                movfpe[.b] ea,rd
bgt[.b] pcrel8                 movtpe[.b] rs,ea_noimm
ble[.w] pcrel16                mulxu sz ea,rd
ble[.b] pcrel8                 neg sz ea
bclr sz imm4,ea_noimm          nop
bclr sz rs,ea_noimm            not sz ea
bnot sz imm4,ea_noimm          or sz ea,rd
bnot sz rs,ea_noimm            orc[.b] imm8,crb
bset sz imm4,ea_noimm          orc[.w] imm16,sr
bset sz rs,ea_noimm            pjmp abs24
bsr[.b] pcrel8                 pjmp @rd
bsr[.w] pcrel16                pjsr abs24
btst sz imm4,ea_noimm          pjsr @rd
btst sz rs,ea_noimm            prtd imm8
clr sz ea                      prtd imm16
cmp[:e][.b] imm8,rd            prts
cmp[:i][.w] imm16,rd           rotl sz ea
cmp[:g].b imm8,ea_noimm        rotr sz ea
cmp[:g][.w] imm16,ea_noimm     rotxl sz ea
Cmp[:g] sz ea,rd               rotxr sz ea
dadd rs,rd                     rtd imm8
divxu sz ea,rd                 rtd imm16
dsub rs,rd                     rts
exts[.b] rd                    scb/f rs,pcrel8
extu[.b] rd                    scb/ne rs,pcrel8
jmp @rd                        scb/eq rs,pcrel8
jmp @(imm8,rd)                 shal sz ea
jmp @(imm16,rd)                shar sz ea
jmp abs16                      shll sz ea
jsr @rd                        shlr sz ea
jsr @(imm8,rd)                 sleep
jsr @(imm16,rd)                stc[.b] crb,ea_noimm
jsr abs16                      stc[.w] sr,ea_noimm
ldm @sp+,(rlist)               stm (rlist),@-sp
link fp,imm8                   sub sz ea,rd
link fp,imm16                  subs sz ea,rd
mov[:e][.b] imm8,rd            subx sz ea,rd
mov[:i][.w] imm16,rd           swap[.b] rd
mov[:l][.w] abs8,rd            tas[.b] ea
mov[:l].b abs8,rd              trapa imm4
mov[:s][.w] rs,abs8            trap/vs
mov[:s].b rs,abs8              tst sz ea
mov[:f][.w] @(disp8,fp),rd     unlk fp
mov[:f][.w] rs,@(disp8,fp)     xch[.w] rs,rd
mov[:f].b @(disp8,fp),rd       xor sz ea,rd
mov[:f].b rs,@(disp8,fp)       xorc.b imm8,crb
mov[:g] sz rs,ea_mem           xorc.w imm16,sr

Chapter 8: Machine Dependent Features 69

     mov[:g] sz ea,rd

8.5 Hitachi SH Dependent Features

8.5.1 Options

as has no additional command-line options for the Hitachi SH family.

8.5.2 Syntax

8.5.2.1 Special Characters

‘!’ is the line comment character.

You can use ‘;’ instead of a newline to separate statements.

Since ‘$’ has no special meaning, you may use it in symbol names.

8.5.2.2 Register Names

You can use the predefined symbols ‘r0’, ‘r1’, ‘r2’, ‘r3’, ‘r4’, ‘r5’, ‘r6’, ‘r7’, ‘r8’, ‘r9’, ‘r10’, ‘r11’, ‘r12’, ‘r13’, ‘r14’, and ‘r15’ to refer to the SH registers.

The SH also has these control registers: pr procedure register (holds return address) pc program counter

mach

macl high and low multiply accumulator registers sr status register gbr global base register vbr vector base register (for interrupt vectors)

8.5.2.3 Addressing Modes

as understands the following addressing modes for the SH. Rn in the following refers to any of the numbered registers, but not the control registers.

Rn Register direct @Rn Register indirect @-Rn Register indirect with pre-decrement @Rn+ Register indirect with post-increment

@(disp, Rn)

Register indirect with displacement

@(R0, Rn) Register indexed

@(disp, GBR)

GBR offset

@(R0, GBR)

GBR indexed

addr

@(disp, PC)

PC relative address (for branch or for addressing memory). The as implementation allows you to use the simpler form addr anywhere a PC relative address is called for; the alternate form is supported for compatibility with other assemblers.

#imm Immediate data

8.5.3 Floating Point

The SH family uses ieee floating-point numbers.

8.5.4 SH Machine Directives

as has no machine-dependent directives for the SH.

8.5.5 Opcodes

For detailed information on the SH machine instruction set, see SH-Microcomputer User’s Manual (Hitachi Micro Systems, Inc.).

as implements all the standard SH opcodes. No additional pseudo-instructions are needed on this family. Note, however, that because as supports a simpler form of PC-relative addressing, you may simply write (for example)

     mov.l  bar,r0

where other assemblers might require an explicit displacement to bar from the program counter:

     mov.l  @(disp, PC)

Here is a summary of SH opcodes:

Legend: Rn a numbered register Rm another numbered register #imm immediate data disp displacement disp8 8-bit displacement disp12 12-bit displacement

add #imm,Rn                    lds.l @Rn+,PR
add Rm,Rn                      mac.w @Rm+,@Rn+
addc Rm,Rn                     mov #imm,Rn
addv Rm,Rn                     mov Rm,Rn
and #imm,R0                    mov.b Rm,@(R0,Rn)
and Rm,Rn                      mov.b Rm,@-Rn
and.b #imm,@(R0,GBR)           mov.b Rm,@Rn
bf disp8                       mov.b @(disp,Rm),R0
bra disp12                     mov.b @(disp,GBR),R0
bsr disp12                     mov.b @(R0,Rm),Rn
bt disp8                       mov.b @Rm+,Rn
clrm                           mov.b @Rm,Rn
clrt                           mov.b R0,@(disp,Rm)
cmp/eq #imm,R0                 mov.b R0,@(disp,GBR)
cmp/eq Rm,Rn                   mov.l Rm,@(disp,Rn)
cmp/ge Rm,Rn                   mov.l Rm,@(R0,Rn)
cmp/gt Rm,Rn                   mov.l Rm,@-Rn
cmp/hi Rm,Rn                   mov.l Rm,@Rn
cmp/hs Rm,Rn                   mov.l @(disp,Rn),Rm
cmp/pl Rn                      mov.l @(disp,GBR),R0
cmp/pz Rn                      mov.l @(disp,PC),Rn
cmp/str Rm,Rn                  mov.l @(R0,Rm),Rn
div0s Rm,Rn                    mov.l @Rm+,Rn
div0u                          mov.l @Rm,Rn
div1 Rm,Rn                     mov.l R0,@(disp,GBR)
exts.b Rm,Rn                   mov.w Rm,@(R0,Rn)
exts.w Rm,Rn                   mov.w Rm,@-Rn
extu.b Rm,Rn                   mov.w Rm,@Rn
extu.w Rm,Rn                   mov.w @(disp,Rm),R0
jmp @Rn                        mov.w @(disp,GBR),R0
jsr @Rn                        mov.w @(disp,PC),Rn
ldc Rn,GBR                     mov.w @(R0,Rm),Rn
ldc Rn,SR                      mov.w @Rm+,Rn
ldc Rn,VBR                     mov.w @Rm,Rn
ldc.l @Rn+,GBR                 mov.w R0,@(disp,Rm)
ldc.l @Rn+,SR                  mov.w R0,@(disp,GBR)
ldc.l @Rn+,VBR                 mova @(disp,PC),R0
lds Rn,MACH                    movt Rn
lds Rn,MACL                    muls Rm,Rn
lds Rn,PR                      mulu Rm,Rn
lds.l @Rn+,MACH                neg Rm,Rn
lds.l @Rn+,MACL                negc Rm,Rn

74 Using as

nop                            stc VBR,Rn
not Rm,Rn                      stc.l GBR,@-Rn
or #imm,R0                     stc.l SR,@-Rn
or Rm,Rn                       stc.l VBR,@-Rn
or.b #imm,@(R0,GBR)            sts MACH,Rn
rotcl Rn                       sts MACL,Rn
rotcr Rn                       sts PR,Rn
rotl Rn                        sts.l MACH,@-Rn
rotr Rn                        sts.l MACL,@-Rn
rte                            sts.l PR,@-Rn
rts                            sub Rm,Rn
sett                           subc Rm,Rn
shal Rn                        subv Rm,Rn
shar Rn                        swap.b Rm,Rn
shll Rn                        swap.w Rm,Rn
shll16 Rn                      tas.b @Rn
shll2 Rn                       trapa #imm
shll8 Rn                       tst #imm,R0
shlr Rn                        tst Rm,Rn
shlr16 Rn                      tst.b #imm,@(R0,GBR)
shlr2 Rn                       xor #imm,R0
shlr8 Rn                       xor Rm,Rn
sleep                          xor.b #imm,@(R0,GBR)
stc GBR,Rn                     xtrct Rm,Rn
stc SR,Rn

8.6 Intel 80960 Dependent Features

8.6.1 i960 Command-line Options

-ACA | -ACA_A | -ACB | -ACC | -AKA | -AKB | -AKC | -AMC

Select the 80960 architecture. Instructions or features not supported by the selected architecture cause fatal errors. ‘-ACA’ is equivalent to ‘-ACA_A’; ‘-AKC’ is equivalent to ‘-AMC’. Synonyms are provided for compatibility with other tools. If none of these options is specified, as will generate code for any instruction or feature that is supported by some version of the 960 (even if this means mixing architectures!). In principle, as will attempt to deduce the minimal sufficient processor type if none is specified; depending on the object code format, the processor type may be recorded in the object file. If it is critical that the as output match a specific architecture, specify that architecture explicitly.

-b Add code to collect information about conditional branches taken, for later optimization using branch prediction bits. (The conditional branch instructions have branch prediction bits in the CA, CB, and CC architectures.) If BR represents a conditional branch instruction, the following represents the code generated by the assembler when ‘-b’ is specified:

call increment routine

             .word   0       # pre-counter
     Label: BR

call increment routine

             .word   0       # post-counter

The counter following a branch records the number of times that branch was not taken; the differenc between the two counters is the number of times the branch was taken. A table of every such Label is also generated, so that the external postprocessor gbr960 (supplied by Intel) can locate all the counters. This table is always labelled ‘__BRANCH_TABLE__’; this is a local symbol to permit collecting statistics for many separate object files. The table is word aligned, and begins with a two-word header. The first word, initialized to 0, is used in maintaining linked lists of branch tables. The second word is a count of the number of entries in the table, which follow immediately: each is a word, pointing to one of the labels illustrated above.

  *NEXT     COUNT: N *BRLAB 1  . . . *BRLAB N

                      BRANCH TABLE layout

The first word of the header is used to locate multiple branch tables, since each object file may contain one. Normally the links are maintained with a call to an initialization routine, placed at the beginning of each function in the file. The GNU C compiler will generate these calls automatically when you give it a ‘-b’ option. For further details, see the documentation of ‘gbr960’.

-norelax Normally, Compare-and-Branch instructions with targets that require displacements greater than 13 bits (or that have external targets) are replaced with the corresponding compare (or ‘chkbit’) and branch instructions. You can use the ‘-norelax’ option to specify that as should generate errors instead, if the target displacement is larger than 13 bits. This option does not affect the Compare-and-Jump instructions; the code emitted for them is always adjusted when necessary (depending on displacement size), regardless of whether you use ‘-norelax’.

8.6.2 Floating Point

as generates ieee floating-point numbers for the directives ‘.float’, ‘.double’, ‘.extended’,

and ‘.single’.

8.6.3 i960 Machine Directives

.bss symbol, length, align Reserve length bytes in the bss section for a local symbol, aligned to the power of two specified by align. length and align must be positive absolute expressions. This directive differs from ‘.lcomm’ only in that it permits you to specify an alignment. See Section 7.29 [.lcomm], page 39.

.extended flonums

.extended expects zero or more flonums, separated by commas; for each flonum, ‘.extended’ emits an ieee extended-format (80-bit) floating-point number.

.leafproc call-lab, bal-lab You can use the ‘.leafproc’ directive in conjunction with the optimized callj instruction to enable faster calls of leaf procedures. If a procedure is known to call no other procedures, you may define an entry point that skips procedure prolog code (and that does not depend on system-supplied saved context), and declare it as the bal-lab using ‘.leafproc’. If the procedure also has an entry point that goes through the normal prolog, you can specify that entry point as call-lab. A ‘.leafproc’ declaration is meant for use in conjunction with the optimized call instruction ‘callj’; the directive records the data needed later to choose between converting the ‘callj’ into a bal or a call. call-lab is optional; if only one argument is present, or if the two arguments are identical, the single argument is assumed to be the bal entry point.

.sysproc name, index

The ‘.sysproc’ directive defines a name for a system procedure. After you define it using ‘.sysproc’, you can use name to refer to the system procedure identified by index when calling procedures with the optimized call instruction ‘callj’. Both arguments are required; index must be between 0 and 31 (inclusive).

8.6.4 i960 Opcodes

All Intel 960 machine instructions are supported; see Section 8.6.1 [i960 Command-line Options], page 76 for a discussion of selecting the instruction subset for a particular 960 architecture.

Some opcodes are processed beyond simply emitting a single corresponding instruction: ‘callj’, and Compare-and-Branch or Compare-and-Jump instructions with target displacements larger than 13 bits.

8.6.4.1 callj

You can write callj to have the assembler or the linker determine the most appropriate form of subroutine call: ‘call’, ‘bal’, or ‘calls’. If the assembly source contains enough information—a ‘.leafproc’ or ‘.sysproc’ directive defining the operand—then as will translate the callj; if not, it will simply emit the callj, leaving it for the linker to resolve.

8.6.4.2 Compare-and-Branch

The 960 architectures provide combined Compare-and-Branch instructions that permit you to store the branch target in the lower 13 bits of the instruction word itself. However, if you specify a branch target far enough away that its address won’t fit in 13 bits, the assembler can either issue an error, or convert your Compare-and-Branch instruction into separate instructions to do the compare and the branch.

Whether as gives an error or expands the instruction depends on two choices you can make: whether you use the ‘-norelax’ option, and whether you use a “Compare and Branch” instruction or a “Compare and Jump” instruction. The “Jump” instructions are always expanded if necessary; the “Branch” instructions are expanded when necessary unless you specify -norelax—in which case as gives an error instead.

These are the Compare-and-Branch instructions, their “Jump” variants, and the instruction pairs they may expand into:

Compare and

Branch  Jump  Expanded to

    bbc             chkbit; bno
    bbs             chkbit; bo
 cmpibe   cmpije    cmpi; be
 cmpibg   cmpijg    cmpi; bg
cmpibge  cmpijge    cmpi; bge
 cmpibl   cmpijl    cmpi; bl
cmpible  cmpijle    cmpi; ble
cmpibno  cmpijno    cmpi; bno
cmpibne  cmpijne    cmpi; bne
 cmpibo   cmpijo    cmpi; bo
 cmpobe   cmpoje    cmpo; be
 cmpobg   cmpojg    cmpo; bg
cmpobge  cmpojge    cmpo; bge
 cmpobl   cmpojl    cmpo; bl
cmpoble  cmpojle    cmpo; ble
cmpobne  cmpojne    cmpo; bne

8.7 M680x0 Dependent Features

8.7.1 M680x0 Options

The Motorola 680x0 version of as has two machine dependent options. One shortens undefined references from 32 to 16 bits, while the other is used to tell as what kind of machine it is assembling for.

You can use the -l option to shorten the size of references to undefined symbols. If the -l option is not given, references to undefined symbols will be a full long (32 bits) wide. (Since as cannot know where these symbols will end up, as can only allocate space for the linker to fill in later. Since as doesn’t know how far away these symbols will be, it allocates as much space as it can.) If this option is given, the references will only be one word wide (16 bits). This may be useful if you want the object file to be as small as possible, and you know that the relevant symbols will be less than 17 bits away.

The 680x0 version of as is most frequently used to assemble programs for the Motorola MC68020 microprocessor. Occasionally it is used to assemble programs for the mostly similar, but slightly different MC68000 or MC68010 microprocessors. You can give as the options ‘-m68000’, ‘-mc68000’, ‘-m68010’, ‘-mc68010’, ‘-m68020’, and ‘-mc68020’ to tell it what processor is the target.

8.7.2 Syntax

The 680x0 version of as uses syntax similar to the Sun assembler. Size modifiers are appended directly to the end of the opcode without an intervening period. For example, write ‘movl’ rather than ‘move.l’.

In the following table apc stands for any of the address registers (‘a0’ through ‘a7’), nothing, (‘’), the Program Counter (‘pc’), or the zero-address relative to the program counter (‘zpc’).

The following addressing modes are understood:

Immediate

‘#digits’

Data Register ‘d0’ through ‘d7’ Address Register

‘a0’ through ‘a7’

Address Register Indirect

‘a0@’ through ‘a7@’

Address Register Postincrement

‘a0@+’ through ‘a7@+’

Address Register Predecrement

‘a0@-’ through ‘a7@-’

Indirect Plus Offset

‘apc@(digits)’

Index ‘apc@(digits,register:size:scale)’

or ‘apc@(register:size:scale)’

Postindex ‘apc@(digits)@(digits,register:size:scale)’

or ‘apc@(digits)@(register:size:scale)’

Preindex ‘apc@(digits,register:size:scale)@(digits)’

or ‘apc@(register:size:scale)@(digits)’

Memory Indirect

‘apc@(digits)@(digits)’

Absolute ‘symbol’, or ‘digits’

For some configurations, especially those where the compiler normally does not prepend an underscore to the names of user variables, the assembler requires a ‘%’ before any use of a register name. This is intended to let the assembler distinguish between user variables and registers named ‘a0’ through ‘a7’, et cetera. The ‘%’ is always accepted, but is only required for some configurations, notably ‘m68k-coff’.

8.7.3 Floating Point

The floating point code is not too well tested, and may have subtle bugs in it.

Packed decimal (P) format floating literals are not supported. Feel free to add the code!

The floating point formats generated by directives are these.

.float Single precision floating point constants. .double Double precision floating point constants.

There is no directive to produce regions of memory holding extended precision numbers, however they can be used as immediate operands to floating-point instructions. Adding a directive to create extended precision numbers would not be hard, but it has not yet seemed necessary.

8.7.4 680x0 Machine Directives

In order to be compatible with the Sun assembler the 680x0 assembler understands the following directives.

.data1 This directive is identical to a .data 1 directive. .data2 This directive is identical to a .data 2 directive. .even This directive is identical to a .align 1 directive. .skip This directive is identical to a .space directive.

8.7.5 Opcodes

8.7.5.1 Branch Improvement

Certain pseudo opcodes are permitted for branch instructions. They expand to the shortest branch instruction that will reach the target. Generally these mnemonics are made by substituting ‘j’ for ‘b’ at the start of a Motorola mnemonic.

The following table summarizes the pseudo-operations. A * flags cases that are more fully described after the table:

               Displacement
               +-------------------------------------------------
               |                68020   68000/10
     Pseudo-Op |BYTE    WORD    LONG    LONG      non-PC relative
               +-------------------------------------------------
          jbsr |bsrs    bsr     bsrl    jsr       jsr
           jra |bras    bra     bral    jmp       jmp

*     jXX |bXXs    bXX     bXXl    bNXs;jmpl bNXs;jmp
*    dbXX |dbXX    dbXX        dbXX; bra; jmpl
*    fjXX |fbXXw   fbXXw   fbXXl             fbNXw;jmp
XX: condition
NX: negative of condition XX

*—see full description below

jbsr

jra These are the simplest jump pseudo-operations; they always map to one particular machine instruction, depending on the displacement to the branch target.

jXX Here, ‘jXX’ stands for an entire family of pseudo-operations, where XX is a conditional branch or condition-code test. The full list of pseudo-ops in this family is:

jhi   jls   jcc   jcs   jne   jeq   jvc
jvs   jpl   jmi   jge   jlt   jgt   jle

For the cases of non-PC relative displacements and long displacements on the 68000 or 68010, as will issue a longer code fragment in terms of NX, the opposite condition to XX. For example, for the non-PC relative case:

   jXX foo

gives

    bNXs oof
    jmp foo
oof:

dbXX The full family of pseudo-operations covered here is

dbhi   dbls   dbcc   dbcs   dbne   dbeq   dbvc
dbvs   dbpl   dbmi   dbge   dblt   dbgt   dble
dbf    dbra   dbt

Other than for word and byte displacements, when the source reads ‘dbXX foo’, as will emit

    dbXX oo1
    bra oo2
oo1:jmpl foo
oo2:

fjXX This family includes

fjne   fjeq   fjge   fjlt   fjgt   fjle   fjf
fjt    fjgl   fjgle  fjnge  fjngl  fjngle fjngt
fjnle  fjnlt  fjoge  fjogl  fjogt  fjole  fjolt
fjor   fjseq  fjsf   fjsne  fjst   fjueq  fjuge
fjugt  fjule  fjult  fjun

For branch targets that are not PC relative, as emits

                    fbNX oof
                    jmp foo
                oof:

when it encounters ‘fjXX foo’.

8.7.5.2 Special Characters

The immediate character is ‘#’ for Sun compatibility. The line-comment character is ‘|’. If a ‘#’ appears at the beginning of a line, it is treated as a comment unless it looks like ‘# line file’, in which case it is treated normally.

8.8 SPARC Dependent Features

8.8.1 Options

The SPARC chip family includes several successive levels (or other variants) of chip, using the same core instruction set, but including a few additional instructions at each level.

By default, as assumes the core instruction set (SPARC v6), but “bumps” the architecture level as needed: it switches to successively higher architectures as it encounters instructions that only exist in the higher levels.

-Av6 | -Av7 | -Av8 | -Asparclite

Use one of the ‘-A’ options to select one of the SPARC architectures explicitly. If you select an architecture explicitly, as reports a fatal error if it encounters an instruction or feature requiring a higher level.

-bump Permit the assembler to “bump” the architecture level as required, but warn whenever it is necessary to switch to another level.

8.8.2 Floating Point

The Sparc uses ieee floating-point numbers.

8.8.3 Sparc Machine Directives

The Sparc version of as supports the following additional machine directives:

.common This must be followed by a symbol name, a positive number, and "bss". This behaves

somewhat like .comm, but the syntax is different.

.half This is functionally identical to .short. .proc This directive is ignored. Any text following it on the same line is also ignored. .reserve This must be followed by a symbol name, a positive number, and "bss". This behaves

somewhat like .lcomm, but the syntax is different.

.seg This must be followed by "text", "data", or "data1". It behaves like .text, .data,

          or .data 1.

86 Using as .skip This is functionally identical to the .space directive. .word On the Sparc, the .word directive produces 32 bit values, instead of the 16 bit values

it produces on many other machines.

8.9 80386 Dependent Features

8.9.1 Options

The 80386 has no machine dependent options.

8.9.2 AT&T Syntax versus Intel Syntax

In order to maintain compatibility with the output of gcc, as supports AT&T System V/386 assembler syntax. This is quite different from Intel syntax. We mention these differences because almost all 80386 documents used only Intel syntax. Notable differences between the two syntaxes are:

• AT&T immediate operands are preceded by ‘$’; Intel immediate operands are undelimited (Intel ‘push 4’ is AT&T ‘pushl $4’). AT&T register operands are preceded by ‘%’; Intel register operands are undelimited. AT&T absolute (as opposed to PC relative) jump/call operands are prefixed by ‘*’; they are undelimited in Intel syntax.

• AT&T and Intel syntax use the opposite order for source and destination operands. Intel ‘add eax, 4’ is ‘addl $4, %eax’. The ‘source, dest’ convention is maintained for compatibility with previous Unix assemblers.

• In AT&T syntax the size of memory operands is determined from the last character of the opcode name. Opcode suffixes of ‘b’, ‘w’, and ‘l’ specify byte (8-bit), word (16-bit), and long (32-bit) memory references. Intel syntax accomplishes this by prefixes memory operands (not the opcodes themselves) with ‘byte ptr’, ‘word ptr’, and ‘dword ptr’. Thus, Intel ‘mov al, byte ptr foo’ is ‘movb foo, %al’ in AT&T syntax.

• Immediate form long jumps and calls are ‘lcall/ljmp $section, $offset’ in AT&T syntax; the Intel syntax is ‘call/jmp far section:offset’. Also, the far return instruction is ‘lret $stack-adjust’ in AT&T syntax; Intel syntax is ‘ret far stack-adjust’.

• The AT&T assembler does not provide support for multiple section programs. Unix style systems expect all programs to be single sections.

8.9.3 Opcode Naming

Opcode names are suffixed with one character modifiers which specify the size of operands. The letters ‘b’, ‘w’, and ‘l’ specify byte, word, and long operands. If no suffix is specified by an instruction and it contains no memory operands then as tries to fill in the missing suffix based on the destination register operand (the last one by convention). Thus, ‘mov %ax, %bx’ is equivalent to ‘movw %ax, %bx’; also, ‘mov $1, %bx’ is equivalent to ‘movw $1, %bx’. Note that this is incompatible with the AT&T Unix assembler which assumes that a missing opcode suffix implies long operand size. (This incompatibility does not affect compiler output since compilers always explicitly specify the opcode suffix.)

Almost all opcodes have the same names in AT&T and Intel format. There are a few exceptions. The sign extend and zero extend instructions need two sizes to specify them. They need a size to sign/zero extend from and a size to zero extend to. This is accomplished by using two opcode suffixes in AT&T syntax. Base names for sign extend and zero extend are ‘movs…’ and ‘movz…’ in AT&T syntax (‘movsx’ and ‘movzx’ in Intel syntax). The opcode suffixes are tacked on to this base name, the from suffix before the to suffix. Thus, ‘movsbl %al, %edx’ is AT&T syntax for “move sign extend from %al to %edx.” Possible suffixes, thus, are ‘bl’ (from byte to long), ‘bw’ (from byte to word), and ‘wl’ (from word to long).

The Intel-syntax conversion instructions

• ‘cbw’ — sign-extend byte in ‘%al’ to word in ‘%ax’,

• ‘cwde’ — sign-extend word in ‘%ax’ to long in ‘%eax’,

• ‘cwd’ — sign-extend word in ‘%ax’ to long in ‘%dx:%ax’,

• ‘cdq’ — sign-extend dword in ‘%eax’ to quad in ‘%edx:%eax’,

are called ‘cbtw’, ‘cwtl’, ‘cwtd’, and ‘cltd’ in AT&T naming. as accepts either naming for these instructions.

Far call/jump instructions are ‘lcall’ and ‘ljmp’ in AT&T syntax, but are ‘call far’ and ‘jump far’ in Intel convention.

8.9.4 Register Naming

Register operands are always prefixes with ‘%’. The 80386 registers consist of

• the 8 32-bit registers ‘%eax’ (the accumulator), ‘%ebx’, ‘%ecx’, ‘%edx’, ‘%edi’, ‘%esi’, ‘%ebp’ (the frame pointer), and ‘%esp’ (the stack pointer).

• the 8 16-bit low-ends of these: ‘%ax’, ‘%bx’, ‘%cx’, ‘%dx’, ‘%di’, ‘%si’, ‘%bp’, and ‘%sp’.

• the 8 8-bit registers: ‘%ah’, ‘%al’, ‘%bh’, ‘%bl’, ‘%ch’, ‘%cl’, ‘%dh’, and ‘%dl’ (These are the high-bytes and low-bytes of ‘%ax’, ‘%bx’, ‘%cx’, and ‘%dx’)

• the 6 section registers ‘%cs’ (code section), ‘%ds’ (data section), ‘%ss’ (stack section), ‘%es’, ‘%fs’, and ‘%gs’.

• the 3 processor control registers ‘%cr0’, ‘%cr2’, and ‘%cr3’.

• the 6 debug registers ‘%db0’, ‘%db1’, ‘%db2’, ‘%db3’, ‘%db6’, and ‘%db7’.

• the 2 test registers ‘%tr6’ and ‘%tr7’.

• the 8 floating point register stack ‘%st’ or equivalently ‘%st(0)’, ‘%st(1)’, ‘%st(2)’, ‘%st(3)’,

  ‘%st(4)’, ‘%st(5)’, ‘%st(6)’, and ‘%st(7)’.

8.9.5 Opcode Prefixes

Opcode prefixes are used to modify the following opcode. They are used to repeat string instructions, to provide section overrides, to perform bus lock operations, and to give operand and address size (16-bit operands are specified in an instruction by prefixing what would normally be 32-bit operands with a “operand size” opcode prefix). Opcode prefixes are usually given as singleline instructions with no operands, and must directly precede the instruction they act upon. For example, the ‘scas’ (scan string) instruction is repeated with:

           repne
           scas

Here is a list of opcode prefixes:

• Section override prefixes ‘cs’, ‘ds’, ‘ss’, ‘es’, ‘fs’, ‘gs’. These are automatically added by specifying using the section:memory-operand form for memory references.

• Operand/Address size prefixes ‘data16’ and ‘addr16’ change 32-bit operands/addresses into 16-bit operands/addresses. Note that 16-bit addressing modes (i.e. 8086 and 80286 addressing modes) are not supported (yet).

• The bus lock prefix ‘lock’ inhibits interrupts during execution of the instruction it precedes. (This is only valid with certain instructions; see a 80386 manual for details).

• The wait for coprocessor prefix ‘wait’ waits for the coprocessor to complete the current instruction. This should never be needed for the 80386/80387 combination.

• The ‘rep’, ‘repe’, and ‘repne’ prefixes are added to string instructions to make them repeat ‘%ecx’ times.

8.9.6 Memory References

An Intel syntax indirect memory reference of the form

section:[base + index*scale + disp]

is translated into the AT&T syntax

section:disp(base, index, scale)

where base and index are the optional 32-bit base and index registers, disp is the optional displacement, and scale, taking the values 1, 2, 4, and 8, multiplies index to calculate the address of the operand. If no scale is specified, scale is taken to be 1. section specifies the optional section register for the memory operand, and may override the default section register (see a 80386 manual for section register defaults). Note that section overrides in AT&T syntax must have be preceded by a ‘%’. If you specify a section override which coincides with the default section register, as will not output any section register override prefixes to assemble the given instruction. Thus, section overrides can be specified to emphasize which section register is used for a given memory operand.

Here are some examples of Intel and AT&T style memory references:

AT&T: ‘-4(%ebp)’, Intel: ‘[ebp - 4]’

base is ‘%ebp’; disp is ‘-4’. section is missing, and the default section is used (‘%ss’ for addressing with ‘%ebp’ as the base register). index, scale are both missing.

AT&T: ‘foo(,%eax,4)’, Intel: ‘[foo + eax*4]’

index is ‘%eax’ (scaled by a scale 4); disp is ‘foo’. All other fields are missing. The section register here defaults to ‘%ds’.

AT&T: ‘foo(,1)’; Intel ‘[foo]’

This uses the value pointed to by ‘foo’ as a memory operand. Note that base and index are both missing, but there is only one ‘,’. This is a syntactic exception.

AT&T: ‘%gs:foo’; Intel ‘gs:foo’

This selects the contents of the variable ‘foo’ with section register section being ‘%gs’.

Absolute (as opposed to PC relative) call and jump operands must be prefixed with ‘*’. If no ‘*’ is specified, as will always choose PC relative addressing for jump/call labels.

Any instruction that has a memory operand must specify its size (byte, word, or long) with an opcode suffix (‘b’, ‘w’, or ‘l’, respectively).

8.9.7 Handling of Jump Instructions

Jump instructions are always optimized to use the smallest possible displacements. This is accomplished by using byte (8-bit) displacement jumps whenever the target is sufficiently close. If a byte displacement is insufficient a long (32-bit) displacement is used. We do not support word (16-bit) displacement jumps (i.e. prefixing the jump instruction with the ‘addr16’ opcode prefix), since the 80386 insists upon masking ‘%eip’ to 16 bits after the word displacement is added.

Note that the ‘jcxz’, ‘jecxz’, ‘loop’, ‘loopz’, ‘loope’, ‘loopnz’ and ‘loopne’ instructions only come in byte displacements, so that it is possible that use of these instructions (gcc does not use them) will cause the assembler to print an error message (and generate incorrect code). The AT&T 80386 assembler tries to get around this problem by expanding ‘jcxz foo’ to

              jcxz cx_zero
              jmp cx_nonzero
     cx_zero: jmp foo
     cx_nonzero:

8.9.8 Floating Point

All 80387 floating point types except packed BCD are supported. (BCD support may be added without much difficulty). These data types are 16-, 32-, and 64- bit integers, and single (32bit), double (64-bit), and extended (80-bit) precision floating point. Each supported type has an opcode suffix and a constructor associated with it. Opcode suffixes specify operand’s data types. Constructors build these data types into memory.

• Floating point constructors are ‘.float’ or ‘.single’, ‘.double’, and ‘.tfloat’ for 32-, 64- , and 80-bit formats. These correspond to opcode suffixes ‘s’, ‘l’, and ‘t’. ‘t’ stands for temporary real, and that the 80387 only supports this format via the ‘fldt’ (load temporary real to stack top) and ‘fstpt’ (store temporary real and pop stack) instructions.

• Integer constructors are ‘.word’, ‘.long’ or ‘.int’, and ‘.quad’ for the 16-, 32-, and 64-bit integer formats. The corresponding opcode suffixes are ‘s’ (single), ‘l’ (long), and ‘q’ (quad). As with the temporary real format the 64-bit ‘q’ format is only present in the ‘fildq’ (load quad integer to stack top) and ‘fistpq’ (store quad integer and pop stack) instructions.

Register to register operations do not require opcode suffixes, so that ‘fst %st, %st(1)’ is equivalent to ‘fstl %st, %st(1)’.

Since the 80387 automatically synchronizes with the 80386 ‘fwait’ instructions are almost never needed (this is not the case for the 80286/80287 and 8086/8087 combinations). Therefore, as suppresses the ‘fwait’ instruction whenever it is implicitly selected by one of the ‘fn…’ instructions. For example, ‘fsave’ and ‘fnsave’ are treated identically. In general, all the ‘fn…’ instructions are made equivalent to ‘f…’ instructions. If ‘fwait’ is desired it must be explicitly coded.

8.9.9 Notes

There is some trickery concerning the ‘mul’ and ‘imul’ instructions that deserves mention. The 16-, 32-, and 64-bit expanding multiplies (base opcode ‘0xf6’; extension 4 for ‘mul’ and 5 for ‘imul’) can be output only in the one operand form. Thus, ‘imul %ebx, %eax’ does not select the expanding multiply; the expanding multiply would clobber the ‘%edx’ register, and this would confuse gcc output. Use ‘imul %ebx’ to get the 64-bit product in ‘%edx:%eax’.

We have added a two operand form of ‘imul’ when the first operand is an immediate mode expression and the second operand is a register. This is just a shorthand, so that, multiplying ‘%eax’ by 69, for example, can be done with ‘imul $69, %eax’ rather than ‘imul $69, %eax, %eax’.

8.10 Z8000 Dependent Features

The Z8000 as supports both members of the Z8000 family: the unsegmented Z8002, with 16 bit addresses, and the segmented Z8001 with 24 bit addresses.

When the assembler is in unsegmented mode (specified with the unsegm directive), an address will take up one word (16 bit) sized register. When the assembler is in segmented mode (specified with the segm directive), a 24-bit address takes up a long (32 bit) register. See Section 8.10.3 [Assembler Directives for the Z8000], page 94, for a list of other Z8000 specific assembler directives.

8.10.1 Options

as has no additional command-line options for the Zilog Z8000 family.

8.10.2 Syntax

8.10.2.1 Special Characters

‘!’ is the line comment character.

You can use ‘;’ instead of a newline to separate statements.

8.10.2.2 Register Names

The Z8000 has sixteen 16 bit registers, numbered 0 to 15. You can refer to different sized groups of registers by register number, with the prefix ‘r’ for 16 bit registers, ‘rr’ for 32 bit registers and ‘rq’ for 64 bit registers. You can also refer to the contents of the first eight (of the sixteen 16 bit registers) by bytes. They are named ‘rnh’ and ‘rnl’. byte registers

     r0l r0h r1h r1l r2h r2l r3h r3l
     r4h r4l r5h r5l r6h r6l r7h r7l

word registers

     r0 r1 r2 r3 r4 r5 r6 r7 r8 r9 r10 r11 r12 r13 r14 r15

long word registers

     rr0 rr2 rr4 rr6 rr8 rr10 rr12 rr14

quad word registers

     rq0 rq4 rq8 rq12

8.10.2.3 Addressing Modes

as understands the following addressing modes for the Z8000:

rn Register direct @rn Indirect register addr Direct: the 16 bit or 24 bit address (depending on whether the assembler is in segmented

or unsegmented mode) of the operand is in the instruction.

address(rn)

Indexed: the 16 or 24 bit address is added to the 16 bit register to produce the final address in memory of the operand.

rn(#imm) Base Address: the 16 or 24 bit register is added to the 16 bit sign extended immediate

displacement to produce the final address in memory of the operand.

rn(rm) Base Index: the 16 or 24 bit register rn is added to the sign extended 16 bit index

register rm to produce the final address in memory of the operand.

#xx Immediate data xx.

8.10.3 Assembler Directives for the Z8000

The Z8000 port of as includes these additional assembler directives, for compatibility with other Z8000 assemblers. As shown, these do not begin with ‘.’ (unlike the ordinary as directives).

segm Generates code for the segmented Z8001. unsegm Generates code for the unsegmented Z8002. name Synonym for .file

global  Synonum for .global

wval Synonym for .word lval Synonym for .long bval Synonym for .byte sval Assemble a string. sval expects one string literal, delimited by single quotes. It

assembles each byte of the string into consecutive addresses. You can use the escape sequence ‘%xx’ (where xx represents a two-digit hexadecimal number) to represent the character whose ascii value is xx. Use this feature to describe single quote and other characters that may not appear in string literals as themselves. For example, the C statement ‘char *a = "he said \"it's 50% off\"";’ is represented in Z8000 assembly language (shown with the assembler output in hex at the left) as

          68652073    sval    'he said %22it%27s 50%25 off%22%00'
          61696420
          22697427
          73203530
          25206F66
          662200
rsect  synonym for .section
block  synonym for .space
even  synonym for .align 1

8.10.4 Opcodes

For detailed information on the Z8000 machine instruction set, see Z8000 Technical Manual.

The following table summarizes the opcodes and their arguments:

rs 16 bit source register rd 16 bit destination register rbs 8 bit source register rbd 8 bit destination register rrs 32 bit source register rrd 32 bit destination register rqs 64 bit source register rqd 64 bit destination register addr 16/24 bit address imm immediate data

adc rd,rs               clrb addr               cpsir @rd,@rs,rr,cc
adcb rbd,rbs            clrb addr(rd)           cpsirb @rd,@rs,rr,cc
add rd,@rs              clrb rbd                dab rbd
add rd,addr             com @rd                 dbjnz rbd,disp7
add rd,addr(rs)         com addr                dec @rd,imm4m1
add rd,imm16            com addr(rd)            dec addr(rd),imm4m1
add rd,rs               com rd                  dec addr,imm4m1
addb rbd,@rs            comb @rd                dec rd,imm4m1

addb rbd,addr           comb addr               decb @rd,imm4m1
addb rbd,addr(rs)       comb addr(rd)           decb addr(rd),imm4m1
addb rbd,imm8           comb rbd                decb addr,imm4m1
addb rbd,rbs            comflg flags            decb rbd,imm4m1
addl rrd,@rs            cp @rd,imm16            di i2
addl rrd,addr           cp addr(rd),imm16       div rrd,@rs
addl rrd,addr(rs)       cp addr,imm16           div rrd,addr
addl rrd,imm32          cp rd,@rs               div rrd,addr(rs)
addl rrd,rrs            cp rd,addr              div rrd,imm16
and rd,@rs              cp rd,addr(rs)          div rrd,rs
and rd,addr             cp rd,imm16             divl rqd,@rs
and rd,addr(rs)         cp rd,rs                divl rqd,addr
and rd,imm16            cpb @rd,imm8            divl rqd,addr(rs)
and rd,rs               cpb addr(rd),imm8       divl rqd,imm32
andb rbd,@rs            cpb addr,imm8           divl rqd,rrs
andb rbd,addr           cpb rbd,@rs             djnz rd,disp7
andb rbd,addr(rs)       cpb rbd,addr            ei i2
andb rbd,imm8           cpb rbd,addr(rs)        ex rd,@rs
andb rbd,rbs            cpb rbd,imm8            ex rd,addr
bit @rd,imm4            cpb rbd,rbs             ex rd,addr(rs)
bit addr(rd),imm4       cpd rd,@rs,rr,cc        ex rd,rs
bit addr,imm4           cpdb rbd,@rs,rr,cc      exb rbd,@rs
bit rd,imm4             cpdr rd,@rs,rr,cc       exb rbd,addr
bit rd,rs               cpdrb rbd,@rs,rr,cc     exb rbd,addr(rs)
bitb @rd,imm4           cpi rd,@rs,rr,cc        exb rbd,rbs
bitb addr(rd),imm4      cpib rbd,@rs,rr,cc      ext0e imm8
bitb addr,imm4          cpir rd,@rs,rr,cc       ext0f imm8
bitb rbd,imm4           cpirb rbd,@rs,rr,cc     ext8e imm8
bitb rbd,rs             cpl rrd,@rs             ext8f imm8
bpt                     cpl rrd,addr            exts rrd
call @rd                cpl rrd,addr(rs)        extsb rd
call addr               cpl rrd,imm32           extsl rqd
call addr(rd)           cpl rrd,rrs             halt
calr disp12             cpsd @rd,@rs,rr,cc      in rd,@rs
clr @rd                 cpsdb @rd,@rs,rr,cc     in rd,imm16
clr addr                cpsdr @rd,@rs,rr,cc     inb rbd,@rs
clr addr(rd)            cpsdrb @rd,@rs,rr,cc    inb rbd,imm16
clr rd                  cpsi @rd,@rs,rr,cc      inc @rd,imm4m1
clrb @rd                cpsib @rd,@rs,rr,cc     inc addr(rd),imm4m1
inc addr,imm4m1         ldb rbd,rs(rx)          mult rrd,addr(rs)
inc rd,imm4m1           ldb rd(imm16),rbs       mult rrd,imm16
incb @rd,imm4m1         ldb rd(rx),rbs          mult rrd,rs
incb addr(rd),imm4m1    ldctl ctrl,rs           multl rqd,@rs
incb addr,imm4m1        ldctl rd,ctrl           multl rqd,addr
incb rbd,imm4m1         ldd @rs,@rd,rr          multl rqd,addr(rs)
ind @rd,@rs,ra          lddb @rs,@rd,rr         multl rqd,imm32
indb @rd,@rs,rba        lddr @rs,@rd,rr         multl rqd,rrs
inib @rd,@rs,ra         lddrb @rs,@rd,rr        neg @rd
inibr @rd,@rs,ra        ldi @rd,@rs,rr          neg addr
iret                    ldib @rd,@rs,rr         neg addr(rd)

jp cc,@rd               ldir @rd,@rs,rr         neg rd
jp cc,addr              ldirb @rd,@rs,rr        negb @rd
jp cc,addr(rd)          ldk rd,imm4             negb addr
jr cc,disp8             ldl @rd,rrs             negb addr(rd)
ld @rd,imm16            ldl addr(rd),rrs        negb rbd
ld @rd,rs               ldl addr,rrs            nop
ld addr(rd),imm16       ldl rd(imm16),rrs       or rd,@rs
ld addr(rd),rs          ldl rd(rx),rrs          or rd,addr
ld addr,imm16           ldl rrd,@rs             or rd,addr(rs)
ld addr,rs              ldl rrd,addr            or rd,imm16
ld rd(imm16),rs         ldl rrd,addr(rs)        or rd,rs
ld rd(rx),rs            ldl rrd,imm32           orb rbd,@rs
ld rd,@rs               ldl rrd,rrs             orb rbd,addr
ld rd,addr              ldl rrd,rs(imm16)       orb rbd,addr(rs)
ld rd,addr(rs)          ldl rrd,rs(rx)          orb rbd,imm8
ld rd,imm16             ldm @rd,rs,n            orb rbd,rbs
ld rd,rs                ldm addr(rd),rs,n       out @rd,rs
ld rd,rs(imm16)         ldm addr,rs,n           out imm16,rs
ld rd,rs(rx)            ldm rd,@rs,n            outb @rd,rbs
lda rd,addr             ldm rd,addr(rs),n       outb imm16,rbs
lda rd,addr(rs)         ldm rd,addr,n           outd @rd,@rs,ra
lda rd,rs(imm16)        ldps @rs                outdb @rd,@rs,rba
lda rd,rs(rx)           ldps addr               outib @rd,@rs,ra
ldar rd,disp16          ldps addr(rs)           outibr @rd,@rs,ra
ldb @rd,imm8            ldr disp16,rs           pop @rd,@rs
ldb @rd,rbs             ldr rd,disp16           pop addr(rd),@rs
ldb addr(rd),imm8       ldrb disp16,rbs         pop addr,@rs
ldb addr(rd),rbs        ldrb rbd,disp16         pop rd,@rs
ldb addr,imm8           ldrl disp16,rrs         popl @rd,@rs
ldb addr,rbs            ldrl rrd,disp16         popl addr(rd),@rs
ldb rbd,@rs             mbit                    popl addr,@rs
ldb rbd,addr            mreq rd                 popl rrd,@rs
ldb rbd,addr(rs)        mres                    push @rd,@rs
ldb rbd,imm8            mset                    push @rd,addr
ldb rbd,rbs             mult rrd,@rs            push @rd,addr(rs)
ldb rbd,rs(imm16)       mult rrd,addr           push @rd,imm16
push @rd,rs             set addr,imm4           subl rrd,imm32
pushl @rd,@rs           set rd,imm4             subl rrd,rrs
pushl @rd,addr          set rd,rs               tcc cc,rd
pushl @rd,addr(rs)      setb @rd,imm4           tccb cc,rbd
pushl @rd,rrs           setb addr(rd),imm4      test @rd
res @rd,imm4            setb addr,imm4          test addr
res addr(rd),imm4       setb rbd,imm4           test addr(rd)
res addr,imm4           setb rbd,rs             test rd
res rd,imm4             setflg imm4             testb @rd
res rd,rs               sinb rbd,imm16          testb addr
resb @rd,imm4           sinb rd,imm16           testb addr(rd)
resb addr(rd),imm4      sind @rd,@rs,ra         testb rbd
resb addr,imm4          sindb @rd,@rs,rba       testl @rd
resb rbd,imm4           sinib @rd,@rs,ra        testl addr

resb rbd,rs             sinibr @rd,@rs,ra       testl addr(rd)
resflg imm4             sla rd,imm8             testl rrd
ret cc                  slab rbd,imm8           trdb @rd,@rs,rba
rl rd,imm1or2           slal rrd,imm8           trdrb @rd,@rs,rba
rlb rbd,imm1or2         sll rd,imm8             trib @rd,@rs,rbr
rlc rd,imm1or2          sllb rbd,imm8           trirb @rd,@rs,rbr
rlcb rbd,imm1or2        slll rrd,imm8           trtdrb @ra,@rb,rbr
rldb rbb,rba            sout imm16,rs           trtib @ra,@rb,rr
rr rd,imm1or2           soutb imm16,rbs         trtirb @ra,@rb,rbr
rrb rbd,imm1or2         soutd @rd,@rs,ra        trtrb @ra,@rb,rbr
rrc rd,imm1or2          soutdb @rd,@rs,rba      tset @rd
rrcb rbd,imm1or2        soutib @rd,@rs,ra       tset addr
rrdb rbb,rba            soutibr @rd,@rs,ra      tset addr(rd)
rsvd36                  sra rd,imm8             tset rd
rsvd38                  srab rbd,imm8           tsetb @rd
rsvd78                  sral rrd,imm8           tsetb addr
rsvd7e                  srl rd,imm8             tsetb addr(rd)
rsvd9d                  srlb rbd,imm8           tsetb rbd
rsvd9f                  srll rrd,imm8           xor rd,@rs
rsvdb9                  sub rd,@rs              xor rd,addr
rsvdbf                  sub rd,addr             xor rd,addr(rs)
sbc rd,rs               sub rd,addr(rs)         xor rd,imm16
sbcb rbd,rbs            sub rd,imm16            xor rd,rs
sc imm8                 sub rd,rs               xorb rbd,@rs
sda rd,rs               subb rbd,@rs            xorb rbd,addr
sdab rbd,rs             subb rbd,addr           xorb rbd,addr(rs)
sdal rrd,rs             subb rbd,addr(rs)       xorb rbd,imm8
sdl rd,rs               subb rbd,imm8           xorb rbd,rbs
sdlb rbd,rs             subb rbd,rbs            xorb rbd,rbs
sdll rrd,rs             subl rrd,@rs
set @rd,imm4            subl rrd,addr
set addr(rd),imm4       subl rrd,addr(rs)

GNU GENERAL PUBLIC LICENSE

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc. 675 Mass Ave, Cambridge, MA 02139, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

1. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

2. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

3. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. You must cause the modified files to carry prominent notices stating that you changed the

files and the date of any change. b. You must cause any work that you distribute or publish, that in whole or in part contains

or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program.

In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

4. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following:

a. Accompany it with the complete corresponding machine-readable source code, which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable.

If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

5. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

6. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

7. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

8. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program.

If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances.

It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

9. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

10. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

11. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

12. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WAR-RANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY APPLICABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPYRIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFOR-MANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DE-FECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

13. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRIT-ING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CON-SEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSI-BILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the program’s name and an idea of what it does.

Copyright (C) 19yy name of author
This program is free software; you can redistribute it and/or
modify it under the terms of the GNU General Public License
as published by the Free Software Foundation; either version 2
of the License, or (at your option) any later version.
This program is distributed in the hope that it will be useful,
but WITHOUT ANY WARRANTY; without even the implied warranty of
MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE.  See the
GNU General Public License for more details.
You should have received a copy of the GNU General Public License
along with this program; if not, write to the Free Software
Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.

Also add information on how to contact you by electronic and paper mail.

If the program is interactive, make it output a short notice like this when it starts in an interactive mode:

Gnomovision version 69, Copyright (C) 19yy name of author
Gnomovision comes with ABSOLUTELY NO WARRANTY; for details
type `show w'.  This is free software, and you are welcome
to redistribute it under certain conditions; type `show c'
for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

     Yoyodyne, Inc., hereby disclaims all copyright
     interest in the program `Gnomovision'
     (which makes passes at compilers) written
     by James Hacker.

signature of Ty Coon, 1 April 1989

     Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

Index 107

Index

#

# . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 #APP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 #NO APP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

$

$ in symbol names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 70

-

-- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 -+ option, VAX/VMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 -a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -A options, i960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 -ad . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -ah . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -al . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -an . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -Asparclite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 -Av6. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 -Av8. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 -b option, i960. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 -D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -D, ignored on VAX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -d, VAX option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -h option, VAX/VMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 -I path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -J, ignored on VAX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -l option, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 -m68000 and related options . . . . . . . . . . . . . . . . . . . . . . . 80 -norelax option, i960. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 -R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 -S, ignored on VAX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -t, ignored on VAX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 -T, ignored on VAX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 -V, redundant on VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 -version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 -W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

.

. (symbol). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 .o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

:

: (label) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

                                \

\" (doublequote character) . . . . . . . . . . . . . . . . . . . . . . . . . 15 \\ (‘\’ character) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 \b (backspace character) . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 \ddd (octal character code) . . . . . . . . . . . . . . . . . . . . . . . . 15 \f (formfeed character) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 \n (newline character) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 \r (carriage return character) . . . . . . . . . . . . . . . . . . . . . . 15 \t (tab) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2

29K support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

A

a.out. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 a.out symbol attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 abort directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 ABORT directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 absent (internal section) . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 absolute section. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 addition, permitted arguments . . . . . . . . . . . . . . . . . . . . . 31 addresses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 addresses, format of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 addressing modes, H8/300 . . . . . . . . . . . . . . . . . . . . . . . . . 57 addressing modes, H8/500 . . . . . . . . . . . . . . . . . . . . . . . . . 64 addressing modes, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . 80 addressing modes, SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 addressing modes, Z8000 . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 advancing location counter . . . . . . . . . . . . . . . . . . . . . . . . . 41 align directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 altered difference tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 AMD 29K floating point (ieee) . . . . . . . . . . . . . . . . . . . . . 55 AMD 29K identifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 AMD 29K line comment character . . . . . . . . . . . . . . . . . 54 AMD 29K line separator . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 AMD 29K machine directives . . . . . . . . . . . . . . . . . . . . . . 55 AMD 29K opcodes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 AMD 29K options (none) . . . . . . . . . . . . . . . . . . . . . . . . . . 54 AMD 29K protected registers . . . . . . . . . . . . . . . . . . . . . . 54 AMD 29K register names . . . . . . . . . . . . . . . . . . . . . . . . . . 54 AMD 29K special purpose registers . . . . . . . . . . . . . . . . 54 AMD 29K statement separator. . . . . . . . . . . . . . . . . . . . . 54 AMD 29K support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 app-file directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 architecture options, i960 . . . . . . . . . . . . . . . . . . . . . . . . . . 76 architecture options, M680x0. . . . . . . . . . . . . . . . . . . . . . . 80 architectures, SPARC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 arguments for addition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 arguments for subtraction . . . . . . . . . . . . . . . . . . . . . . . . . . 31 arguments in expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 arithmetic functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 arithmetic operands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 as version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ascii directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 asciz directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 assembler internal logic error . . . . . . . . . . . . . . . . . . . . . . . 22 assembler, and linker. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 assembly listings, enabling . . . . . . . . . . . . . . . . . . . . . . . . . . 7 assigning values to symbols . . . . . . . . . . . . . . . . . . . . . 25, 36 attributes, symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 auxiliary attributes, COFF symbols . . . . . . . . . . . . . . . . 28 auxiliary symbol information, COFF . . . . . . . . . . . . . . . 35 Av7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

B

backslash (\\) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 backspace (\b) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 bignum/flonum (internal section). . . . . . . . . . . . . . . . . . . 22 bignums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 binary integers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 bitfields, not supported on VAX. . . . . . . . . . . . . . . . . . . . 53 block. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 block directive, AMD 29K . . . . . . . . . . . . . . . . . . . . . . . . . 55 branch improvement, M680x0 . . . . . . . . . . . . . . . . . . . . . . 82 branch improvement, VAX . . . . . . . . . . . . . . . . . . . . . . . . . 51 branch recording, i960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 branch statistics table, i960 . . . . . . . . . . . . . . . . . . . . . . . . 76 bss directive, i960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 bss section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21, 24 bus lock prefixes, i386. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 bval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 byte directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

C

call instructions, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 callj, i960 pseudo-opcode . . . . . . . . . . . . . . . . . . . . . . . . . 78 carriage return (\r) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 character constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 character escape codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 character, single . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 characters used in symbols . . . . . . . . . . . . . . . . . . . . . . . . . 13 COFF auxiliary symbol information. . . . . . . . . . . . . . . . 35 COFF named section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 COFF structure debugging. . . . . . . . . . . . . . . . . . . . . . . . . 45 COFF symbol attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 COFF symbol descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 COFF symbol storage class . . . . . . . . . . . . . . . . . . . . . . . . 43 COFF symbol type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 COFF symbols, debugging . . . . . . . . . . . . . . . . . . . . . . . . . 35 COFF value attribute. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 comm directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 command line conventions . . . . . . . . . . . . . . . . . . . . . . . . . . 4 command-line options ignored, VAX . . . . . . . . . . . . . . . 49 comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 comments, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 comments, removed by preprocessor . . . . . . . . . . . . . . . . 11 common directive, SPARC . . . . . . . . . . . . . . . . . . . . . . . . . . 85 common variable storage . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 compare and jump expansions, i960 . . . . . . . . . . . . . . . . 79 compare/branch instructions, i960. . . . . . . . . . . . . . . . . . 78 conditional assembly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 constant, single character . . . . . . . . . . . . . . . . . . . . . . . . . . 15 constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 constants, bignum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 constants, character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 constants, converted by preprocessor . . . . . . . . . . . . . . . 11 constants, floating point . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 constants, integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 constants, number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 constants, string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 continuing statements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 conversion instructions, i386 . . . . . . . . . . . . . . . . . . . . . . . 88 coprocessor wait, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 cputype directive, AMD 29K . . . . . . . . . . . . . . . . . . . . . . 55 current address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 current address, advancing . . . . . . . . . . . . . . . . . . . . . . . . . 41

D

data and text sections, joining . . . . . . . . . . . . . . . . . . . . . . 9 data directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 data section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 data1 directive, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 data2 directive, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 debuggers, and symbol order . . . . . . . . . . . . . . . . . . . . . . . 25 debugging COFF symbols. . . . . . . . . . . . . . . . . . . . . . . . . . 35 decimal integers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 def directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 deprecated directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 desc directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 descriptor, of a.out symbol . . . . . . . . . . . . . . . . . . . . . . . . 28 dfloat directive, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 difference (internal section) . . . . . . . . . . . . . . . . . . . . . . . . 22 difference tables altered . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 difference tables, warning . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 dim directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 directives and instructions . . . . . . . . . . . . . . . . . . . . . . . . . 13 directives, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 directives, machine independent . . . . . . . . . . . . . . . . . . . . 33 directives, Z8000 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 displacement sizing character, VAX . . . . . . . . . . . . . . . . 53 dot (symbol). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 double directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 double directive, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 double directive, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . 82 double directive, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 doublequote (\"). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

E

eight-byte integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 eject directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 else directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 empty expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 endef directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 endif directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 EOF, newline must precede . . . . . . . . . . . . . . . . . . . . . . . . 13 equ directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 error messsages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 escape codes, character . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 even. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 even directive, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 expression arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 expressions, empty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 expressions, integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 extended directive, i960. . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 extern directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

F

faster processing (-f) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ffloat directive, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 file directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 file directive, AMD 29K . . . . . . . . . . . . . . . . . . . . . . . . . . 55 file name, logical . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34, 37 files, including . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 files, input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 fill directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 filling memory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 float directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 float directive, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 float directive, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 float directive, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 floating point numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 floating point numbers (double) . . . . . . . . . . . . . . . . . . . . 36 floating point numbers (single) . . . . . . . . . . . . . . . . . 37, 44 floating point, AMD 29K (ieee) . . . . . . . . . . . . . . . . . . . . 55 floating point, H8/300 (ieee) . . . . . . . . . . . . . . . . . . . . . . . 58 floating point, H8/500 (ieee) . . . . . . . . . . . . . . . . . . . . . . . 64 floating point, i386. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 floating point, i960 (ieee). . . . . . . . . . . . . . . . . . . . . . . . . . . 77 floating point, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 floating point, SH (ieee). . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 floating point, SPARC (ieee) . . . . . . . . . . . . . . . . . . . . . . . 85 floating point, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 flonums. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 format of error messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 format of warning messages . . . . . . . . . . . . . . . . . . . . . . . . . 6 formfeed (\f) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 functions, in expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 fwait instruction, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

G

gbr960, i960 postprocessor . . . . . . . . . . . . . . . . . . . . . . . . . 76 gfloat directive, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 global . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 global directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 grouping data. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

H

H8/300 addressing modes . . . . . . . . . . . . . . . . . . . . . . . . . . 57 H8/300 floating point (ieee) . . . . . . . . . . . . . . . . . . . . . . . . 58 H8/300 line comment character . . . . . . . . . . . . . . . . . . . . 57 H8/300 line separator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 H8/300 machine directives (none) . . . . . . . . . . . . . . . . . . 58 H8/300 opcode summary . . . . . . . . . . . . . . . . . . . . . . . . . . 58 H8/300 options (none) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 H8/300 registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 H8/300 size suffixes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 H8/300 support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 H8/500 addressing modes . . . . . . . . . . . . . . . . . . . . . . . . . . 64 H8/500 floating point (ieee) . . . . . . . . . . . . . . . . . . . . . . . . 64 H8/500 line comment character . . . . . . . . . . . . . . . . . . . . 63 H8/500 line separator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 H8/500 machine directives (none) . . . . . . . . . . . . . . . . . . 64 H8/500 opcode summary . . . . . . . . . . . . . . . . . . . . . . . . . . 65 H8/500 options (none) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 H8/500 registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 H8/500 support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 half directive, SPARC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 hexadecimal integers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 hfloat directive, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 hword directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

I

i386 conversion instructions . . . . . . . . . . . . . . . . . . . . . . . . 88 i386 floating point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 i386 fwait instruction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 i386 immediate operands. . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i386 jump optimization . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 i386 jump, call, return . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i386 jump/call operands . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i386 memory references . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 i386 mul, imul instructions . . . . . . . . . . . . . . . . . . . . . . . . . 92 i386 opcode naming. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i386 opcode prefixes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 i386 options (none) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i386 register operands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i386 registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 i386 sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i386 size suffixes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i386 source, destination operands. . . . . . . . . . . . . . . . . . . 87 i386 support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i386 syntax compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i80306 support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 i960 architecture options . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 i960 branch recording . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 i960 callj pseudo-opcode. . . . . . . . . . . . . . . . . . . . . . . . . . 78 i960 compare and jump expansions . . . . . . . . . . . . . . . . . 79 i960 compare/branch instructions . . . . . . . . . . . . . . . . . . 78 i960 floating point (ieee) . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 i960 machine directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 i960 opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 i960 options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 i960 support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 ident directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 identifiers, AMD 29K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 if directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 ifdef directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 ifndef directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 ifnotdef directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 immediate character, M680x0 . . . . . . . . . . . . . . . . . . . . . . 84 immediate character, VAX . . . . . . . . . . . . . . . . . . . . . . . . . 53 immediate operands, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . 87 imul instruction, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 include directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 include directive search path . . . . . . . . . . . . . . . . . . . . . . . 8 indirect character, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 infix operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 inhibiting interrupts, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . 89 input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 input file linenumbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 instruction set, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 instruction summary, H8/300 . . . . . . . . . . . . . . . . . . . . . . 58 instruction summary, H8/500 . . . . . . . . . . . . . . . . . . . . . . 65 instruction summary, SH. . . . . . . . . . . . . . . . . . . . . . . . . . . 71 instruction summary, Z8000 . . . . . . . . . . . . . . . . . . . . . . . . 95 instructions and directives . . . . . . . . . . . . . . . . . . . . . . . . . 13 int directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 int directive, H8/300 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 int directive, H8/500 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 int directive, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 int directive, SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 integer expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 integer, 16-byte . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 integer, 8-byte . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 integers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 integers, 16-bit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 integers, 32-bit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 integers, binary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 integers, decimal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 integers, hexadecimal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 integers, octal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 integers, one byte . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 internal as sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 invocation summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

J

joining text and data sections . . . . . . . . . . . . . . . . . . . . . . . 9 jump instructions, i386. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 jump optimization, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 jump/call operands, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

L

label (:) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 labels. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 lcomm directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 ld . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 leafproc directive, i960. . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 length of symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 lflags directive (ignored) . . . . . . . . . . . . . . . . . . . . . . . . . 39 line comment character . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 line comment character, AMD 29K. . . . . . . . . . . . . . . . . 54 line comment character, H8/300. . . . . . . . . . . . . . . . . . . . 57 line comment character, H8/500. . . . . . . . . . . . . . . . . . . . 63 line comment character, M680x0 . . . . . . . . . . . . . . . . . . . 84 line comment character, SH . . . . . . . . . . . . . . . . . . . . . . . . 70 line comment character, Z8000 . . . . . . . . . . . . . . . . . . . . . 93 line directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 line directive, AMD 29K . . . . . . . . . . . . . . . . . . . . . . . . . . 55 line numbers, in input files . . . . . . . . . . . . . . . . . . . . . . . . . . 5 line numbers, in warnings/errors . . . . . . . . . . . . . . . . . . . . 6 line separator character . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 line separator, AMD 29K . . . . . . . . . . . . . . . . . . . . . . . . . . 54 line separator, H8/300 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 line separator, H8/500 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 line separator, SH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 line separator, Z8000. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 lines starting with # . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 linker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 linker, and assembler. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 list directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 listing control, turning off . . . . . . . . . . . . . . . . . . . . . . . . . . 41 listing control, turning on . . . . . . . . . . . . . . . . . . . . . . . . . . 40 listing control: new page . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 listing control: paper size . . . . . . . . . . . . . . . . . . . . . . . . . . 42 listing control: subtitle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 listing control: title line . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 listings, enabling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ln directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 local common symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 local labels, retaining in output . . . . . . . . . . . . . . . . . . . . . 8 local symbol names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 location counter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 location counter, advancing . . . . . . . . . . . . . . . . . . . . . . . . 41 logical file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34, 37 logical line number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 logical line numbers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 long directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 long directive, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 lval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

M

M680x0 addressing modes. . . . . . . . . . . . . . . . . . . . . . . . . . 80 M680x0 architecture options . . . . . . . . . . . . . . . . . . . . . . . 80 M680x0 branch improvement. . . . . . . . . . . . . . . . . . . . . . . 82 M680x0 directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 M680x0 floating point. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 M680x0 immediate character. . . . . . . . . . . . . . . . . . . . . . . 84 M680x0 line comment character . . . . . . . . . . . . . . . . . . . . 84 M680x0 opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 M680x0 options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 M680x0 pseudo-opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 M680x0 size modifiers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 M680x0 support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 M680x0 syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 machine dependencies. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 machine directives, AMD 29K. . . . . . . . . . . . . . . . . . . . . . 55 machine directives, H8/300 (none) . . . . . . . . . . . . . . . . . 58 machine directives, H8/500 (none) . . . . . . . . . . . . . . . . . 64 machine directives, i960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 machine directives, SH (none) . . . . . . . . . . . . . . . . . . . . . . 71 machine directives, SPARC . . . . . . . . . . . . . . . . . . . . . . . . 85 machine directives, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 machine independent directives . . . . . . . . . . . . . . . . . . . . 33 machine instructions (not covered) . . . . . . . . . . . . . . . . . . 3 machine-independent syntax . . . . . . . . . . . . . . . . . . . . . . . 11 manual, structure and purpose . . . . . . . . . . . . . . . . . . . . . . 2 memory references, i386. . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 merging text and data sections . . . . . . . . . . . . . . . . . . . . . . 9 messages from as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 minus, permitted arguments . . . . . . . . . . . . . . . . . . . . . . . 31 mnemonics for opcodes, VAX . . . . . . . . . . . . . . . . . . . . . . 51 mnemonics, H8/300. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 mnemonics, H8/500. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 mnemonics, SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 mnemonics, Z8000 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 mul instruction, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 multi-line statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

N

name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 named section (COFF) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 named sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 names, symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 naming object file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 new page, in listings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 newline (\n) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 newline, required at file end . . . . . . . . . . . . . . . . . . . . . . . . 13 nolist directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 null-terminated strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 number constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 numbered subsections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 numbers, 16-bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 numeric values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

O

object file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 object file format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 object file name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 obsolescent directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 octa directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 octal character code (\ddd) . . . . . . . . . . . . . . . . . . . . . . . . 15 octal integers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 opcode mnemonics, VAX. . . . . . . . . . . . . . . . . . . . . . . . . . . 51 opcode naming, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 opcode prefixes, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 opcode suffixes, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 opcode summary, H8/300 . . . . . . . . . . . . . . . . . . . . . . . . . . 58 opcode summary, H8/500 . . . . . . . . . . . . . . . . . . . . . . . . . . 65 opcode summary, SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 opcode summary, Z8000 . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 opcodes for AMD 29K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 opcodes, i960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 opcodes, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 operand delimiters, i386. . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 operand notation, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 operands in expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 operator precedence. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 operators, in expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 operators, permitted arguments . . . . . . . . . . . . . . . . . . . . 30 option summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 options for AMD29K (none) . . . . . . . . . . . . . . . . . . . . . . . 54 options for i386 (none) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 options for SPARC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 options for VAX/VMS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 options, all versions of as . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 options, command line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 options, H8/300 (none) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 options, H8/500 (none) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 options, i960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 options, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 options, SH (none). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 options, Z8000 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 org directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 other attribute, of a.out symbol . . . . . . . . . . . . . . . . . . . 28 output file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

P

padding the location counter . . . . . . . . . . . . . . . . . . . . . . . 33 page, in listings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 paper size, for listings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 pass1 (internal section) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 paths for .include. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 patterns, writing in memory . . . . . . . . . . . . . . . . . . . . . . . 37 plus, permitted arguments . . . . . . . . . . . . . . . . . . . . . . . . . 31 precedence of operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 precision, floating point . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 prefix operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 prefixes, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 preprocessing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 preprocessing, turning on and off . . . . . . . . . . . . . . . . . . . 11 primary attributes, COFF symbols . . . . . . . . . . . . . . . . . 28 proc directive, SPARC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 protected registers, AMD 29K . . . . . . . . . . . . . . . . . . . . . 54 pseudo-opcodes, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 pseudo-ops for branch, VAX . . . . . . . . . . . . . . . . . . . . . . . 51 pseudo-ops, machine independent . . . . . . . . . . . . . . . . . . 33 psize directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 purpose of gnu as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Q

quad directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 quad directive, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

R

register names, AMD 29K. . . . . . . . . . . . . . . . . . . . . . . . . . 54 register names, VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 register operands, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 registers, H8/300 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 registers, H8/500 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 registers, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 registers, SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 registers, Z8000 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 relocation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 relocation example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 repeat prefixes, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 reserve directive, SPARC . . . . . . . . . . . . . . . . . . . . . . . . . 85 return instructions, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 rsect. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

S

sbttl directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 scl directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 search path for .include. . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 sect directive, AMD 29K . . . . . . . . . . . . . . . . . . . . . . . . . . 55 section directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 section override prefixes, i386 . . . . . . . . . . . . . . . . . . . . . . 89 section-relative addressing . . . . . . . . . . . . . . . . . . . . . . . . . 20 sections. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 sections in messages, internal. . . . . . . . . . . . . . . . . . . . . . . 22 sections, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 sections, named . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 seg directive, SPARC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 segm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 set directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 SH addressing modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 SH floating point (ieee) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 SH line comment character. . . . . . . . . . . . . . . . . . . . . . . . . 70 SH line separator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 SH machine directives (none) . . . . . . . . . . . . . . . . . . . . . . 71 SH opcode summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 SH options (none) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 SH registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 SH support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 short directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 single character constant . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 single directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 single directive, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 sixteen bit integers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 sixteen byte integer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 size directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 size modifiers, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 size prefixes, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 size suffixes, H8/300 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 sizes operands, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 skip directive, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 skip directive, SPARC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 source program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 source, destination operands; i386 . . . . . . . . . . . . . . . . . . 87 space directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 SPARC architectures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 SPARC floating point (ieee). . . . . . . . . . . . . . . . . . . . . . . . 85 SPARC machine directives . . . . . . . . . . . . . . . . . . . . . . . . . 85 SPARC options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 SPARC support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 special characters, M680x0 . . . . . . . . . . . . . . . . . . . . . . . . . 84 special purpose registers, AMD 29K . . . . . . . . . . . . . . . . 54 stabd directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 stabn directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 stabs directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 stabx directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 standard as sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 standard input, as input file . . . . . . . . . . . . . . . . . . . . . . . . . 4 statement on multiple lines . . . . . . . . . . . . . . . . . . . . . . . . 13 statement separator character . . . . . . . . . . . . . . . . . . . . . . 13 statement separator, AMD 29K . . . . . . . . . . . . . . . . . . . . 54 statement separator, H8/300 . . . . . . . . . . . . . . . . . . . . . . . 57 statement separator, H8/500 . . . . . . . . . . . . . . . . . . . . . . . 63 statement separator, SH . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 statement separator, Z8000 . . . . . . . . . . . . . . . . . . . . . . . . 93 statements, structure of . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 stopping the assembly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 string constants. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 string literals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 structure debugging, COFF . . . . . . . . . . . . . . . . . . . . . . . . 45 subexpressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 subtitles for listings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 subtraction, permitted arguments . . . . . . . . . . . . . . . . . . 31 summary of options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 supporting files, including . . . . . . . . . . . . . . . . . . . . . . . . . . 39 suppressing warnings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 sval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 symbol attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 symbol attributes, a.out . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 symbol attributes, COFF . . . . . . . . . . . . . . . . . . . . . . . . . . 28 symbol descriptor, COFF . . . . . . . . . . . . . . . . . . . . . . . . . . 35 symbol names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 symbol names, ‘$’ in . . . . . . . . . . . . . . . . . . . . . . . . . . . 63, 70 symbol names, local . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 symbol names, temporary . . . . . . . . . . . . . . . . . . . . . . . . . . 26 symbol storage class (COFF) . . . . . . . . . . . . . . . . . . . . . . 43 symbol type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 symbol type, COFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 symbol value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 symbol value, setting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 symbol values, assigning . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 symbol, common. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 symbol, making visible to linker . . . . . . . . . . . . . . . . . . . . 38 symbolic debuggers, information for . . . . . . . . . . . . . . . . 44 symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 symbols with lowercase, VAX/VMS . . . . . . . . . . . . . . . . 50 symbols, assigning values to . . . . . . . . . . . . . . . . . . . . . . . . 36 symbols, local common . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 syntax compatibility, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . 87 syntax, M680x0. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 syntax, machine-independent. . . . . . . . . . . . . . . . . . . . . . . 11 sysproc directive, i960 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

T

tab (\t) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 tag directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 temporary symbol names . . . . . . . . . . . . . . . . . . . . . . . . . . 26 text and data sections, joining . . . . . . . . . . . . . . . . . . . . . . 9 text directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 text section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 tfloat directive, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 title directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 trusted compiler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 turning preprocessing on and off . . . . . . . . . . . . . . . . . . . 11 type directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 type of a symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

U

undefined section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 unsegm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 use directive, AMD 29K . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

V

val directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 value attribute, COFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 value of a symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 VAX bitfields not supported . . . . . . . . . . . . . . . . . . . . . . . 53 VAX branch improvement . . . . . . . . . . . . . . . . . . . . . . . . . 51 VAX command-line options ignored . . . . . . . . . . . . . . . . 49 VAX displacement sizing character . . . . . . . . . . . . . . . . . 53 VAX floating point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 VAX immediate character. . . . . . . . . . . . . . . . . . . . . . . . . . 53 VAX indirect character . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 VAX machine directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 VAX opcode mnemonics . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 VAX operand notation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 VAX register names. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 VAX support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Vax-11 C compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 VAX/VMS options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 version of as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 VMS (VAX) options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

W

warning for altered difference tables . . . . . . . . . . . . . . . . . 8 warning messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 warnings, suppressing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 whitespace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 whitespace, removed by preprocessor . . . . . . . . . . . . . . . 11 wide floating point directives, VAX. . . . . . . . . . . . . . . . . 50 word directive. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 word directive, H8/300 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 word directive, H8/500 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 word directive, i386 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 word directive, SH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 word directive, SPARC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 writing patterns in memory . . . . . . . . . . . . . . . . . . . . . . . . 37 wval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

Z

Z800 addressing modes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 Z8000 directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 Z8000 line comment character. . . . . . . . . . . . . . . . . . . . . . 93 Z8000 line separator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 Z8000 opcode summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 Z8000 options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 Z8000 registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 Z8000 support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 zero-terminated strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

116 Using as

Table of Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Structure of this Manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 as, the GNU Assembler. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.3 Object File Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.4 Command Line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.5 Input Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.6 Output (Object) File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.7 Error and Warning Messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

2 Command-Line Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1 Enable Listings: -a[dhlns] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.2 -D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.3 Work Faster: -f. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4 .include search path: -I path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.5 Difference Tables: -K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.6 Include Local Labels: -L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.7 Name the Object File: -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.8 Join Data and Text Sections: -R. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.9 Announce Version: -v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.10 Suppress Warnings: -W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3 Syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3.1 Pre-Processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.2 Whitespace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.3 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.4 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.5 Statements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.6 Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

3.6.1 Character Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.6.1.1 Strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.6.1.2 Characters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

3.6.2 Number Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 3.6.2.1 Integers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.6.2.2 Bignums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.6.2.3 Flonums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ii Using as

4 Sections and Relocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

4.1 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 4.2 ld Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 4.3 as Internal Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 4.4 Sub-Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.5 bss Section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

5 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

5.1 Labels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.2 Giving Symbols Other Values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.3 Symbol Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 5.4 The Special Dot Symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 5.5 Symbol Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

5.5.1 Value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 5.5.2 Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 5.5.3 Symbol Attributes: a.out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

5.5.3.1 Descriptor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.5.3.2 Other. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

5.5.4 Symbol Attributes for COFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.5.4.1 Primary Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 5.5.4.2 Auxiliary Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

6 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

6.1 Empty Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 6.2 Integer Expressions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

6.2.1 Arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 6.2.2 Operators. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 6.2.3 Prefix Operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 6.2.4 Infix Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

7 Assembler Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

7.1 .abort. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 7.2 .ABORT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 7.3 .align abs-expr , abs-expr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 7.4 .app-file string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 7.5 .ascii "string". . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 7.6 .asciz "string". . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 7.7 .byte expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 7.8 .comm symbol , length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 7.9 .data subsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.10 .def name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.11 .desc symbol, abs-expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.12 .dim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.13 .double flonums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.14 .eject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.15 .else. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.16 .endef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.17 .endif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.18 .equ symbol, expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 7.19 .extern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.20 .file string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.21 .fill repeat , size , value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.22 .float flonums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.23 .global symbol, .globl symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 7.24 .hword expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 7.25 .ident . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 7.26 .if absolute expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 7.27 .include "file" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 7.28 .int expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 7.29 .lcomm symbol , length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 7.30 .lflags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 7.31 .line line-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.32 .ln line-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.33 .list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.34 .long expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.35 .nolist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 7.36 .octa bignums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 7.37 .org new-lc , fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 7.38 .psize lines , columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 7.39 .quad bignums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 7.40 .sbttl "subheading" . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 7.41 .scl class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 7.42 .section name, subsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 7.43 .set symbol, expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 7.44 .short expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 7.45 .single flonums . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 7.46 .size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 7.47 .space size , fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 7.48 .stabd, .stabn, .stabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 7.49 .tag structname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 7.50 .text subsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 7.51 .title "heading". . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 7.52 .type int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 7.53 .val addr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 7.54 .word expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

7.55 Deprecated Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

8 Machine Dependent Features . . . . . . . . . . . . . . . . . . . . . . 49

8.1 VAX Dependent Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 8.1.1 VAX Command-Line Options . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 8.1.2 VAX Floating Point. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 8.1.3 Vax Machine Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 8.1.4 VAX Opcodes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 8.1.5 VAX Branch Improvement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 8.1.6 VAX Operands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 8.1.7 Not Supported on VAX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

8.2 AMD 29K Dependent Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 8.2.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 8.2.2 Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

8.2.2.1 Special Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 8.2.2.2 Register Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

8.2.3 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 8.2.4 AMD 29K Machine Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 8.2.5 Opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

8.3 H8/300 Dependent Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 8.3.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 8.3.2 Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

8.3.2.1 Special Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 8.3.2.2 Register Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 8.3.2.3 Addressing Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57

8.3.3 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 8.3.4 H8/300 Machine Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 8.3.5 Opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

8.4 H8/500 Dependent Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 8.4.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 8.4.2 Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

8.4.2.1 Special Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 8.4.2.2 Register Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 8.4.2.3 Addressing Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

8.4.3 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 8.4.4 H8/500 Machine Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 8.4.5 Opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

8.5 Hitachi SH Dependent Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 8.5.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 8.5.2 Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

8.5.2.1 Special Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 8.5.2.2 Register Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

8.5.2.3 Addressing Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

8.5.3 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 8.5.4 SH Machine Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 8.5.5 Opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

8.6 Intel 80960 Dependent Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 8.6.1 i960 Command-line Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 8.6.2 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 8.6.3 i960 Machine Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 8.6.4 i960 Opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

8.6.4.1 callj . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 8.6.4.2 Compare-and-Branch . . . . . . . . . . . . . . . . . . . . . . . . . . 78

8.7 M680x0 Dependent Features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 8.7.1 M680x0 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 8.7.2 Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 8.7.3 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 8.7.4 680x0 Machine Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 8.7.5 Opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

8.7.5.1 Branch Improvement. . . . . . . . . . . . . . . . . . . . . . . . . . . 82 8.7.5.2 Special Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

8.8 SPARC Dependent Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 8.8.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 8.8.2 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 8.8.3 Sparc Machine Directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

8.9 80386 Dependent Features. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 8.9.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 8.9.2 AT&T Syntax versus Intel Syntax . . . . . . . . . . . . . . . . . . . . . . . 87 8.9.3 Opcode Naming. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 8.9.4 Register Naming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 8.9.5 Opcode Prefixes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 8.9.6 Memory References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 8.9.7 Handling of Jump Instructions . . . . . . . . . . . . . . . . . . . . . . . . . . 91 8.9.8 Floating Point . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 8.9.9 Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92

8.10 Z8000 Dependent Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 8.10.1 Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 8.10.2 Syntax. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

8.10.2.1 Special Characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 8.10.2.2 Register Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 8.10.2.3 Addressing Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

8.10.3 Assembler Directives for the Z8000 . . . . . . . . . . . . . . . . . . . . . 94 8.10.4 Opcodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 vi Using as

GNU GENERAL PUBLIC LICENSE. . . . . . . . . . . . . . . . . 99

Preamble. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND

MODIFICATION. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 How to Apply These Terms to Your New Programs . . . . . . . . . . . . . . . . . . . 105

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107